### v1

# WarmSwap: Sharing Dependencies for Accelerating Cold Starts in Serverless Functions

Abstract
This work presents WarmSwap, a novel provider-side cold-start optimization for serverless computing. This optimization reduces cold-start time when booting and loading dependencies at runtime inside a function container. Previous approaches to the optimization of cold starts tend to fall into two categories: optimizing the infrastructure of serverless computing to benefit all serverless functions; or function-specific tuning for individual serverless functions. In contrast, WarmSwap offers a broad middle ground, which optimizes entire categories of serverless functions. WarmSwap eliminates the need to initialize middleware or software dependencies when launching a new serverless container, by migrating a pre-initialized live dependency image to the new function instance. WarmSwap respects the provider’s cache constraints, as a single pre-warmed dependency image in the cache is shared among all serverless functions requiring that software dependency image. WarmSwap has been tested on seven representative functions from FunctionBench. The functions are chosen to compare with previous work. In those tests, WarmSwap accelerates cold-start executions for those serverless functions with large dependency requirements by a factor ranging from 1.2 to 2.2.
1 Introduction
Serverless computing is a flexible and popular cloud computing methodology that provides users with a simpler and cheaper way to deploy applications. It has quickly gained attention across various cloud providers, including AWS [3], Azure [19], Google[11], and Alibaba [2]. For the user, serverless computing offers robust auto-scaling and an economical pay-as-you-go model, in contrast to reserving an instance for a fixed time. Users only need to submit their code as a serverless function in the cloud and later trigger the function’s execution based on a request or event. Finally, the user’s code is executed in instances (e.g., a virtual machine or a Docker container) provided by cloud providers.
However, many of the financial benefits are undermined by the cold-start problem. In serverless computing, the cloud provider loads user function into the serverless instance only upon user invocation, forming the basis of the “pay-as-you-go” pricing model. Typically, the cloud retains the serverless instance in memory for a certain period (i.e., during a keep-alive time).
If an invocation occurs when no instance is alive in the system, the function experiences a cold start, which is significantly slower due to instance initialization. If the next invocation happens within the keep-alive time, the function experiences a warm start, during which the provider does not need to load and initialize a new instance. Cold starts can also occur during spikes in invocations when the system requires additional instances of a given serverless function. However, as reported by Azure [22], this affects only a small percentage of applications (less than 1%).
Given the significance of the cold-start problem [10], recent years have witnessed a surge in related research efforts from both academia and industry [6, 8, 21, 23, 26, 20, 15, 24]. To address the overhead of cold start, these researchers propose various system designs that leverage techniques such as caching, remote memory, and checkpoint/restore. Their ultimate goal is to achieve a faster cold start by accelerating or bypassing certain initialization phases.
Optimizations for serverless cold start can be classified into function-specific tuning and infrastructure optimization based on their targets. Function-specific tuning focuses on individual functions, such as creating and caching pre-initialized images [23, 8, 6], generating function seeds [26], and recording and pre-fetching hot pages [25]. Infrastructure optimization typically targets the entire system or shareable resources to benefit different functions. Examples include building a more efficient storage or caching system [5, 1], maintaining pre-initialized Zygote containers [21, 16], and sharing idle containers across different functions [14].
Function-specific tuning typically results in very fast cold starts because it can manage every detail of the initialization phase for a specific function. However, implementing these optimizations for all serverless functions on the cloud can incur substantial costs for the cloud provider. In contrast, infrastructure optimizations have lower costs for large-scale cloud providers because they naturally benefit all serverless functions running on the cloud. Furthermore, infrastructure optimizations are easier to implement in real-world cloud environments because they only require involvement from the cloud providers.
Figure 1: The relationship of invocation rate $\lambda$ versus number of expected cold starts with keep-alive time $T=15$ minutes and the function distribution coming from Azure [22]. Each serverless function was assigned to a bucket $[x,x+0.001]$, where the $\lambda$ of the serverless function was between $x$ and $x+0.001$ invocations per minute. The y-value for a bucket is the number of functions that fall into the given bucket. The y-values for each bucket were normalized so that the sum of the values for the buckets adds to one.
In practice, production cloud providers tend to emphasize more infrastructure optimization, due to the skewed workload in the real world, as illustrated in Figure 1. Function-specific tuning is suitable for functions with invocation rates around $1/T$, where a significant number of expected cold starts occur to each function. However, according to an Azure Function dataset [22], 50% of functions experience fewer than $1.4$ cold starts per day, assuming a keep-alive time of 15 minutes. Therefore, it is costly and unnecessary to maintain pre-initialized images or record hot pages in RAM for all serverless functions.
Although some infrastructure optimization for serverless container caching and launching have been successfully implemented in production [1, 5], there remains a lack of infrastructure optimization for loading software dependencies. For instance, AWS Lambda can launch cold starts for function containers up to 10 GB in size within 50 ms [5], but it still initializes software dependencies from scratch (see experiments in Section 2.3). The time required to load software dependencies within the container can sometimes exceed the execution time [21]. Recently, this initialization time has been exacerbated by the development of AI and large language models, as AI-related functions often require larger and more complex dependencies.
The existing work targets reducing dependency-loading latency, but deploying those solutions in real-world serverless systems still presents challenges. SOCK [21] and FlashCube [16] have proposed new container frameworks that aim to bypass container launch and dependency-loading phases by preserving Zygote containers and container components on serverless worker nodes. However, preserving pre-initialized containers or their components on every serverless worker node, or remotely migrating the entire container, is costly. It also restricts user flexibility to use their own custom containers, a feature commonly supported by large serverless providers like AWS [3].
Finally, this work is complementary to previous work, and can effectively coexist with previous approaches. The function-specific tuning like catalyzer [8] or vHive [25] optimize the frequently used serverless functions (i.e. around $1/T$ in Figure 1). This work optimizes the “long tail” of the distribution of buckets for serverless functions (see caption of Figure 1), as documented by Azure [22]. And finally, infrastructure optimizations benefit [5, 1] all serverless functions.
The WarmSwap Approach:
In this paper, we present WarmSwap, targeting the dependency loading overhead during cold starts. In WarmSwap, the provider maintains a pool of live dependency images in RAM. A live dependency image is an image stored in memory that represents a checkpoint of the software dependencies. Typically, these images contain popular middleware combinations, such as Python with NumPy, Python with Torch, or Python with NumPy and Torch.
WarmSwap is designed to benefit serverless providers by also optimizing the massive number of less frequently invoked functions (the “long tail” of the distribution), which is usually not practical in a function-specific approach. WarmSwap brings several benefits. First, providers only need to maintain a limited number of live dependency images in memory, which can be shared by a wide variety of serverless functions that utilize that middleware. The pre-warming cost for the provider depends solely on the number of live dependency images, not on the number of serverless functions employed by multiple users. However, the benefit of pre-initialization is shared among all serverless functions using the middleware. Second, WarmSwap respects isolation for serverless functions. This is because the provider primarily transfers pages from public dependency images, which do not include user source code.
By using WarmSwap, the provider can accelerate the cold start by “pre-warming” a popular dependency pool in either the local worker nodes or a remote central node. When creating new serverless instances for cold starts, WarmSwap can first check whether a necessary software dependency already exists in the dependency image pool. If so, WarmSwap conducts a live migration of the required dependency image from the pool to new serverless containers. This saves time otherwise spent on booting runtimes and loading software dependencies.
Furthermore, the WarmSwap approach is practical because it focuses on widely used, public software dependencies. A study [21] reveals that 36% of “import” in 876,000 Python projects are from just 20 packages. Therefore, providers can concentrate on maintaining a relatively small number of dependency images across numerous serverless functions used by all users.
An experimental implementation of WarmSwap was evaluated by running offline AWS Lambda containers on AWS EC2 nodes. In our tests, loading libraries such as “sklearn+pandas”, “numpy+keras”, and “numpy+torch” through WarmSwap was up to 2.2x, 3.2x, and 2.5x faster, respectively, compared to traditional AWS Lambda startup processes. Furthermore, WarmSwap can save 85.4% of memory space compared to the function-specific approach [23] when warming up 10 different serverless functions sharing the same dependency. Therefore, for functions sharing a common dependency, WarmSwap represents a new type of infrastructure optimization using pre-warmed dependency images, offering several advantages over previous approaches in the literature.
We summarize our contributions as follows.
  1. 1.
    This paper proposes WarmSwap, a new infrastructure optimization targeting the loading of a software dependency to accelerate serverless cold starts, benefiting the many less frequently invoked serverless functions as well.
  2. 2.
    WarmSwap is a cost-effective solution for dependency-loading, as the provider stores middleware only once in a separate dependency image, independent of the entire container. A single pre-warmed dependency image can be shared between different serverless functions and worker nodes.
  3. 3.
    This work is targeted to benefit the real-world distributions of serverless function invocations with their “long tail”, as documented by an Azure study (see Figure 1).
The rest of this paper is organized as: Section 2 presents the background and motivation of our paper; Section 3 presents the design and implementation details of WarmSwap; In Section  4, we evaluate the performance of WarmSwap; We then discuss related work in Section 5, and conclude in Section 6.
2 Background and Motivation
This section introduces the background for serverless computing and the issue of cold starts in Section 2.1. Section 2.2 motivates the need for WarmSwap by analyzing the relationship between cold starts and function invocation patterns in Azure. Section 2.3 then discusses an existing performance bottleneck in a commercial cloud platform such as AWS Lambda.
2.1 Cold Start in Serverless Computing
This section introduces the background of container-based serverless computing from a cold-start perspective. Figure 2 shows the general process of such a cold start.
Figure 2: The process of serverless cold starts in the current production cloud: Different optimization methods may involve one or more steps in the process.
In Figure 2, while the specifics may differ across platforms, most serverless architectures typically consist of two core components: a Worker Manager and a Resource Orchestration unit. In response to a function invocation from the front end, a request for capacity is sent to the Worker Manager, which acts as a stateful load balancer. The Worker Manager continuously monitors the capacity of each unique function on the platform. If sufficient capacity exists (e.g., idle function instances are available), the Worker Manager directs the front end to route the request to an idle instance. Conversely, if resources are unavailable (triggering a cold start), it instructs Resource Orchestration to provision a new function instance with adequate CPU and RAM.
Recall also that the user has uploaded the container function to the provider’s Container Registry. Next, the provider will download the user container from the Container Registry and start it. The container initiates the serverless runtime to boot the entire serverless function. Then, the serverless runtime starts a specific function runtime (e.g., Python), and the program loads the necessary software dependencies (e.g., NumPy) and the user-specific handler code. At this point, the new function instance is ready to handle user requests.
To minimize the cold-start overhead, one needs to skip or expedite one or more steps in the cold-start optimization. When designing a function-specific tuning for the cold start, one can consider all the steps specific to a serverless function. Previous works have designed end-to-end, function-specific tuning for targeted serverless functions and achieved substantial reductions in cold-start latency. However, when performing infrastructure optimization for the entire system, one can only optimize steps that are identical or similar across all serverless functions, such as container storage and loading.
In this paper, we argue that runtime booting and dependency-loading qualify for infrastructure optimization because they are common processes that can benefit multiple serverless functions. Thus, we propose WarmSwap as a novel infrastructure optimization aimed at reducing the time required for runtime booting and dependency-loading.
2.2 Impact of Invocation Patterns on Cold Starts
In this section, we aim to address two questions:
  1. 1)
    How frequently does a cold start occur for a serverless function?
  2. 2)
    How does a skewed workload impact the optimization choices of the cloud provider?
To simplify the discussion, we focus on situations where a cold start occurs due to the absence of any active instances, as this scenario accounts for more than $99\%$ of cold starts in production systems [22].
To address question 1, we model the occurrence of cold starts using an exponential distribution. The frequency of cold starts for a serverless function depends on its invocation pattern. For extreme cases, a function that is never invoked or invoked very frequently (e.g., every second) would not experience cold starts. Let $\lambda$ denote the invocation rate (calls per minute), and $T$ denote the keep-alive time. The probability of no invocation occurring during the keep-alive time follows an exponential distribution:
$P_{no\_inv}=e^{-\lambda T}$ (1)
Here, $P_{no\_inv}$ represents the probability that no invocation happens during the keep-alive timeout, leading to a potential cold start.
The number of expected cold starts that occur within $D$ minutes can be expressed as:
$E_{cs}(\lambda)=D\lambda P_{no\_inv}=D\lambda e^{-\lambda T}$ (2)
which has a maximum when $dE/d\lambda=0,\lambda=1/T$.
Figure 1 shows the relation of $E_{cs}$ versus $\lambda$ when $T=10\;min$ and $D=1440\;min$ (a day). Only functions with invocation rates around $1/T$ have a large number of expected cold starts, functions that are less or more frequently invoked have fewer cold starts. Cloud providers like Azure [19] and AWS [3] configure dynamic keep-alive time from $5$ to $30$ minutes. The different keep-alive time changes the position of the maximum point, but the shapes are similar.
To answer question two, we draw the density histogram of the actual function distribution in the same Figure 1. For function-specific tuning, assuming the cost of optimizing a function is $c$ and the benefits of an optimized cold start are $w$. Functions that are worth function-specific tuning must satisfy the following:
$wE_{cs}(\lambda)>c$ (3)
However, from the histogram, the serverless workload in production can be extremely skewed. Over $50\%$ of functions have very a low invocation rate (less than 0.001 calls per minute) and $E_{cs}$, which experience less than $1.4$ times of cold starts. Thus, those functions are hard to qualify equation 3, which means function-specific tuning is not economic for them.
For those functions whose invocation rate cannot qualify equation 3, only infrastructure optimization can help to reduce their cold-start overhead. One successful optimization example from the provider is AWS Lambda, which published its optimization of function creation and container loading in a recent paper [5]. The design of AWS lambda assumes that serverless instances are highly cacheable, in which most scaling is driven by a smaller number of images, and the images are based on common base layers. With a three-tier caching system and container deduplication, AWS can cold start a 10 GB serverless container within 50 ms. Thus, the loading and booting of the container are lightweight in current commercial serverless systems.
In this paper, we hope WarmSwap can further extend the infrastructure optimization for the benefit of the many serverless functions that can not apply function-specific tuning.
2.3 Performance Analysis of the Cold Start for Current AWS Lambda Functions
This section analyzes the performance of cold-start overhead in a production cloud on AWS [1, 5]. The goal is to analyze which components of the current AWS Lambda architecture still have room for optimization. None of the experiments in this section refer to WarmSwap.
For analysis, we built serverless function containers to create AWS Lambda functions, and then invoke them by a script running on an AWS EC2 node. All Lambda functions are configured with a 10 GB memory limitation to maximize the CPU resources. To collect cold-start times, the script needs to wait until the end of keep-alive time between invocations of the same function. There are several timers in the analysis script and the function containers to track the time of each cold-start process. The entire time for the cold start is calculated from the script that sends the request to the serverless function and receives the result.
Note that the caching system of AWS Lambda would bring a bias to a naive experimental design. For example, if the user uploads a new function container, the first cold start will be extremely slow as compared to the following ones. Moreover, if the functions are frequently invoked, AWS upgrades the function to a different cache that will improve its cold starts. (See “AWS three-tier caching” in [5].) The improved caching scheme continues to be used even when the function is deleted and then recreated. In order to achieve reproducible numerical results, we upload a container and then invoke the serverless function many times, until the cold-start results are stable.
Serverless Functions Container Size (MB) Major Dependencies
helloworld 180.00 None
json_dumps_load 180.00 urllib, json
pyase 180.11 pyaes, string
chameleon 180.42 chameleon
lr_serving 379.52 sklearn, pandas
cnn_serving 1386.73 numpy, keras
rnn_serving 5602.46 numpy, torch
Table 1: Serverless functions adopted from FunctionBench [12]
Table 1 presents serverless functions from a representative serverless benchmark suite called FunctionBench [12, 13]. We have chosen FunctionBench, for ease of comparison with previous work [25]. All these functions are written in Python and they use some representative Python packages like “numpy”, “torch”, “keras”, and “sklearn”. Table 1 lists both packages and the size of specific function container images. Installing large packages also significantly increases the image size of function containers. Section 4.1 describes more details of these functions.
Figure 3: Cold-start latency breakdown for serverless functions in a traditional AWS Lambda function (not WarmSwap).
Figure 3 shows the breakdown of intermediate times for cold starts in this analysis. The entire cold-start time can be broken down into to following parts:
  • Network: The process of network communications between invoker and user functions.
  • Container: The process of creating the AWS containers.
  • Boot: The process of booting Python and the AWS Lambda interface [4].
  • Dependency Initialization: The process of importing required middleware, referred to as serverless software dependencies.
  • Execution: The process of execution of user handler.
From Figure 3, the dependency initialization dominates the cold start of AWS Lambda functions. In “lr_serving”, “cnn_serving”, and “rnn_serving”, the dependency initialization occupies $43\%$, $74\%$ and $68\%$ of the cold-start overhead, respectively. The time for the dependency initialization component is even higher than that for the execution component for the three AWS Lambda functions. The dependency initialization component is a major problem for serverless functions in production. This problem occurs frequently because the packages they need are so common.
Note that the time for the container component is almost constant (about $0.5~{}s$), even for containers that have different image sizes. This shows that AWS has already optimized the time for the container component, by improving the AWS caching and file systems [5, 1].
In summary, the time of the dependency initialization component often dominates the time for the other components. Therefore, improving the dependency initialization component overhead is important both to serverless providers and to users of the current production cloud. This motivates a novel method focusing on the dependency initialization overhead, as seen in Section 3.
(a) System Architecture
(b) Setup Phase
(c) Runtime Phase
Figure 4: System architecture and workflow: (a) Architectureal overview of WarmSwap and its dependency pool. (b) In the setup phase, the user registers a serverless function, and WarmSwap generates the corresponding image in the dependency pool. (c) In the runtime phase, pages are copied from the image between provider worker nodes and user function containers.
3 WarmSwap Design
This section introduces WarmSwap, a prototype system for enabling “live dependency image” on serverless platforms. WarmSwap is motivated by the goal of improving the cold start by skipping the dependency loading phase. WarmSwap enables a serverless function container to seamlessly migrate the dependency image of its pre-initialized software dependencies from a provider’s worker node. For example, assume there is a serverless function based on Python and numpy. The function can use pre-initialized memory pages of Python and numpy through migration. This eliminates the otherwise large runtime overhead of first initializing Python and numpy.
WarmSwap leverages the power of a shared dependency pool, enabling significant performance gains for providers. Even though individual functions differ, multiple users whose functions have a common dependency can benefit from the same pre-initialized dependency image from the pool. The cloud provider can then support such a dependency pool for all serverless users. In general, the pool contains several pre-initialized dependency images for popular runtimes and packages (e.g., Python + numpy, Python + torch, etc.).
We first provide an overview of the system’s architecture in Section 3.1, highlighting the key components and their interactions. Next, we detail the implementation of each WarmSwap component, including the generation of the pool of dependency images in Section 3.2 and the design choices of migration in Section 3.3. Finally, Section 3.4 describes the implementation.
3.1 Design Overview
WarmSwap is designed to reduce dependency loading overhead for serverless functions. The key idea is to migrate pre-initialized dependencies instead of starting them from scratch. Here, migrate means to copy a process from the host (which could be a local or remote node) to the user container over the network. The system design is based on the characteristics of such dependencies. A container for a serverless function usually contains both user-specific code and a middleware platform with its own package dependencies. We refer to this middleware as a serverless software dependency. These containers generally share three characteristics, described below, which determine the design and performance of the WarmSwap workflow.
Commonality:
Although there are numerous serverless software dependencies, and applications can use different combinations of dependencies, many applications are based on a small number of popular packages (e.g., NumPy, sklearn, torch). Providing migration services for those popular packages can benefit complex serverless functions with large dependencies.
Sparsity:
Most applications do not need all functions from a package, which means they only need access to small parts of the dependency memory pages. In this situation, it is unnecessary to load all the memory pages before the execution of user functions.
Lightweight:
Keeping runtime and software dependencies in the memory is much more lightweight than keeping the whole container. Thus, the provider can afford to maintain a pool of dependency images on their work node.
Figure 4a describes the components of WarmSwap on the provider work node and function container. On the provider worker node, a Dependency Manager orchestrates the management of in-memory or "live" dependency images, which are images stored in memory that represent a checkpoint of pre-initialized software dependencies (e.g., Python + numpy). It is the central hub for communication with user function containers, facilitating migration services. The Dependency Manager maintains metadata about existing live dependency images and manages the memory pages associated with them. When a migration request is initiated by a serverless container, the Dependency Manager accepts the request, transfers the essential metadata required for process restoration, and provides page transfer services to the serverless function container.
In the container for the user function, a Migration Client orchestrates the migration of memory pages on the user side. Upon startup of the container, it initiates a migration request for the Dependency Manager. Upon receiving confirmation, it starts the Dependency Loader to orchestrate the restoration of the runtime and the loading of dependency images, guided by the provided metadata.
Rather than loading all memory pages immediately, the Dependency Loader restores the memory pages of the dependency by lazy migration. The Dependency Loader restores the dependency without loading the actual memory pages; instead, it initiates a Page Server to handle this task later. When the user functions execute and access memory pages of the dependency, it triggers page faults, and the Page Server handles them. This Page Server communicates with the dependency pool to efficiently load specific pages whenever required by the process.
3.2 Generation of Dependency Images
Figure 4b shows the setup phase of WarmSwap, which dumps a serverless software dependency into a dependency image in the dependency pool. The serverless software dependency process (refer to the target process in the following paragraphs) in this example is a Python script importing numpy and then executing the user handler. We initiate the dumping of the process after numpy has been loaded but before the import and execution of the user code (the user handler in Figure 4a). This dump timing ensures that the migrated process can support different user handlers and execute them immediately after migration.
The target process is divided into two distinct components during the dumping process: process metadata, preserved within Dependency Metadata; and hot memory pages, residing securely within the Dependency pool. The process metadata encompasses essential information for restoring the process framework, including file descriptors, memory segment sizes, and other relevant configuration details. The dependency pool preserves the full memory pages of the target process within its memory domain. To enable efficient page transfer, the Dependency Manager creates a dedicated page server process on the source for each target process. The source page server uses the network to communicate with page servers on the destination that need the target process.
To further accelerate the process generation and optimize resource utilization, we employ checkpoint images on disk. When dumping target processes, we create checkpoint images in order on disk to further optimize the generation process. With images on disk available, we do not need to execute the Python script from the beginning. Instead, we can directly restore the checkpoint and then create live dependency images. This is useful for those large and less popular target processes because the provider can save costs by putting them on disk instead of memory.
WarmSwap requires only a modest pool of dependency images, because of the commonality and lightweight characteristics of serverless software dependencies. Thus, cloud providers only undertake limited costs associated with dumping and maintaining dependent images.
3.3 Migration of Dependency Images
Figure 4c describes the runtime details of initializing a WarmSwap container. Due to the sparsity of the memory pages required by a dependency image, we leverage a strategy known as lazy migration. This approach dynamically transfers pages only as needed, significantly reducing overhead and enhancing performance.
The migration client orchestrates the migration process, which has three distinct steps.
The first step of migration is to transport process metadata to the destination (i.e., to the serverless function container). After receiving the migration request from the migration client, the dependency manager looks up the dependency metadata for the requested dependency image. If the dependency image exists, the dependency manager transports process metadata to the dependency loader.
The second step of migration is to start a page server, which handles the “Page Faults” of the dependency to be loaded by using the “userfaultfd” feature of Linux. The page server accesses the dependency pool through the network, to load those pages that actually are needed by the serverless function.
In the third step, the dependency loader then restores the serverless software dependencies based on the process metadata. The dependency loader: (a) restores the memory structure; and then (b) reconnects file descriptors in the new environment. After migration, the new process continues to import and execute the user handler.
However, the dependency loader does not load all of the memory pages in the dependency image. Instead, it starts a page server that connects to a specific dependency image inside the dependency pool on the worker node. Thus, when the process accesses specific memory pages, it triggers a “Page Fault”, which the page server handles using “userfaultfd”.
For handling the “Page Fault”, WarmSwap could implement either of two possible design choices:
Lazy restore: When handling the “Page Fault”, only the pages required by the user functions are transferred. In this case, the Page Server needs to access the dependency pool every time a page fault occurs. Thus, the execution of the user function is slowed down by the number of pages the function needs. However, this method guarantees the transfer of only the minimum number of pages.
Bulk restore: When handling the first “Page Fault”, the Page Server first transfers the required pages. And then, in the background, it loads the remaining pages in parallel until another page fault occurs. In this case, the Page Server needs to access the dependency pool only once. Therefore, the execution will not be impacted by further page faults and remote communication. Although transferring all pages increases communication overhead, this should not impact the user functions because the Page Server can work in parallel with the user functions.
WarmSwap chooses bulk restore because it generally benefits all functions without impacting the warm starts (see Section 4.4). In contrast, lazy restore only works well in fewer page fault cases. In a third possibility, one might record the subset of pages that the serverless functions need and only transfer them. For example, previous work [25] records and pre-fetch hot pages for each target serverless function. However, recall that this method involves function-specific tuning, which does not fulfill the requirements of WarmSwap.
3.4 Implementation
We implement a prototype of WarmSwap with AWS Lambda containers and CRIU [7] (Checkpoint and Restore In Userspace). WarmSwap uses AWS Lambda container templates (Version Python 3.9) to build serverless functions and collect performance results. The dependency loader is embedded into the boot process of an AWS Lambda container. When a serverless container receives invocation, it first starts the dependency loader instead of the original “bootstrap” script. Then the dependency loader loads dependency images into this container and continually executes the AWS Lambda Runtime Interface [4] and the user code.
Note that WarmSwap is a provider-side implementation. It was necessary to simulate WarmSwap on an AWS EC2 node instead of the online AWS Lambda environment. Nevertheless, all experiments employ a standard AWS Lambda container and simply execute the container with the EC2 environment. The EC2 environment was required because AWS Lambda does not allow the user to control the placement of serverless functions on nodes. In particular, it cannot guarantee that a serverless function will be placed on the same node as its associated pool of dependency images.
The prototype modifies CRIU (Version 3.9) to implement live migration of dependency images. We add two new features supporting the design described in Section 3.2 and Section 3.3.
First, we modified the “criu dump” command to support the dependency pool. This ensures that the “criu dump” command will always keep the live dependency image in memory (not killed after migration) and managed by the dependency manager of WarmSwap. Additionally, we extend the live dependency image to accept multiple requests.
Second, we modified the “criu pageserver” command to support the lazy restore policy described in Section 3.3. Note that there is a “--lazy page” parameter of the “criu pageserver” command. However, this parameter migrates all pages upon the execution start. We need to implement WarmSwap’s lazy restore policy: loading pages one-by-one. Similarly, CRIU’s --lazy page parameter does not fully support WarmSwap’s bulk restore policy. Some smaller modifications were also required for this policy.
Finally, it is important that the Python runtime and packages are installed in the same path and with the same version on both the worker node and the container. Otherwise, the “criu restore” cannot find the necessary files and reconnect the file descriptors.
4 Evaluation
We first present the experimental setup in Section 4.1, including the experimental platform and benchmarks for serverless functions. The implementation was described in Section 3.4. The evaluation answers the following research questions:
  • RQ1: How well does WarmSwap improve the cold-start performance of different serverless functions? (Section 4.2)
  • RQ2: How does each component in WarmSwap contribute to the cold start of different serverless functions? (Section 4.3)
  • RQ3: How well does live migration in WarmSwap improve the cold-start execution compared with standard migration using simple checkpoint images? (Section 4.4)
  • RQ4: How do different migration policies (lazy restore and bulk restore) work in WarmSwap? (Section 4.4)
  • RQ5: How well does WarmSwap perform when sharing dependency images across multiple serverless functions? (Section 4.5)
4.1 Experimental Setup
We use a “r5b.large” AWS EC2 machine with two vCPUs, 16 GB memory, and AWS EBS 200 GB SSD, to conduct our experiments. The O/S is Amazon Linux with Linux kernel 5.10.199, and the gcc version is 10.5.0. As described in Section 3.4, an AWS EC2 node is required to simulate a provider-side WarmSwap implementation, but standard AWS Lambda server containers are used in all experiments. Moreover, all serverless containers are built based on the “public.ecr.aws/lambda/python:3.9” template. The serverless containers are started with “--privileged” to allow the execution of CRIU.
All experiments use Python-based functions from the FunctionBench [12, 13] serverless benchmark suite. Recall that the middleware packages needed by these functions are described in Table 1 in Section 2.3. The purposes of these serverless functions are as follows:
  • helloworld: Minimal function with no dependency packages.
  • pyase: Text encryption with an AES block-cipher.
  • chameleon: HTML table rendering.
  • json_dumps_load: Serialization and de-serialization of JSON.
  • lr_serving: Serving a logistic regressions model.
  • cnn_serving: Serving an image classification model based on CNN.
  • rnn_serving: Serving a sequence generation model based on RNN.
Functions needing model parameters (namely, lr _serving, cnn _serving, rnn _serving) download their parameters from AWS S3 when executing cold starts. The dependency images created for each serverless function are based on Python and Python packages described in Table 1.
The cold-start latency discussed in this section refers to the times for: booting at runtime; loading dependencies; and executing the user functions, unless otherwise specified. To record the cold-start latency, we first start the serverless container in the worker node and wait for the invocations to eliminate the overhead of container creation.
4.2 Comparing WarmSwap and Baseline
To answer RQ1, we evaluate the cold-start latency of WarmSwap and the Baseline with all functions described in Section 4.1. Baseline refers to a traditional AWS Lambda container running offline. The Baseline runs the same containers that were built for online AWS Lambda in Section 2.3 on an EC2 machine. In this experiment, WarmSwap uses bulk restore (see Section 3.3) because it has the best performance. These serverless containers directly boot the AWS Lambda Runtime Interface during cold starts and use a Python environment and packages pre-installed inside the serverless container.
(a) Cold Start
(b) Warm Start
Figure 5: Comparison of WarmSwap versus Baseline. The software dependencies for each function are specified in Table 1 of Section 2.3. Note that lr_serving, cnn_serving, and rnn_serving all have much larger dependency images. For this reason, WarmSwap is shown to be much faster on cold starts than the Baseline in these cases, by a ratio of 1.2, 1.8, and 2.2. Further, there is almost no degradation of performance during warm starts.
Figure 5a shows that WarmSwap outperforms the Baseline on cold-start latency for serverless functions that require large dependencies. WarmSwap is slightly slower than Baseline for the lightweight serverless functions (namely, helloworld, pyase, chameleon, and json_ dumps_load). However, WarmSwap significantly outperforms Baseline for serverless functions that need larger packages (i.e., “pandas”, “numpy”, “scikit-learn”, “torch”, and “keras”). WarmSwap is 1.2x, 1.8x, and 2.2x faster than Baseline for “lr _serving”, “cnn _serving”, and “rnn _serving”, respectively.
Figure 5b shows the warm-start latency of WarmSwap and Baseline. For all six serverless functions, WarmSwap has similar warm-start latencies with Baseline.
The cold start results show that WarmSwap has a performance advantage only on certain serverless functions. Specifically, WarmSwap is designed to solve the large dependency-loading problem in the serverless cold start. The lightweight functions only import standard and small Python packages. So directly booting those functions is faster than using WarmSwap. The lightweight functions incur additional overhead because WarmSwap needs to communicate with the dependency manager and migrate the dependency image into the containers, which increases some overhead.
The warm start results show that WarmSwap’s performance is almost the same as that of the Baseline containers. During cold starts, the user code is already executed and necessary pages are already installed. WarmSwap does not add overhead to the warm start of the same function. Furthermore, there is no difference between the execution times for regular software dependencies and for migrated dependencies.
4.3 Components of Cold Start Execution Time
To further analyze WarmSwap when executing a cold start, and answer RQ2, we decompose the cold-start time of WarmSwap and Baseline into three parts. Recall the architecture of WarmSwap from Figure 4a. Communication is the process by which the “migration client” communicates with the dependency manager and receives the process metadata. The process metadata is a special data structure associated with checkpoints that support CRIU’s lazy migration protocol (see Section 3.2). Migration is the process in which the dependency loader executes until the start of user-specific handler functions. The definitions of Execution, Boot, and Depedency Initialization are the same as those in Section 2.3.
Figure 6: Cold-start latency breakdown for WarmSwap (H) and Baseline (B). The performance gain of WarmSwap arises because the migration is faster than the software dependency initialization of a Baseline container.
Figure 6 shows the breakdown of cold-start latency for WarmSwap. The communication time is almost constant for all tested serverless functions, ranging from 25 ms to 94 ms according to the sizes of process metadata. Table 3 shows the sizes of process metadata for dependency images for each function. The size of process metadata is much smaller than the checkpoint image of the entire dependency. Since WarmSwap already migrates most of the memory by the page server, it does not need to transfer much additional information in the process metadata.
Compared with the dependency-loading time of Baseline, WarmSwap has a faster migration time in the case of functions that have larger dependencies. WarmSwap is 2.2x, 3.2x, and 2.5x faster than Baseline for booting dependencies for “lr _serving”, “cnn _serving” and “rnn _serving”, respectively. That explains how WarmSwap helps with the cold-start time. Since all memory pages are pre-initialized in the dependency image, user functions directly use them after migration. Within WarmSwap, the migration time dominates the cold start because migration is the key phase that transfers the dependency image that is needed by the user function.
4.4 Analysis of Benefits of WarmSwap Components
This section evaluates the contribution of the different optimization techniques used by WarmSwap in cold starts: live migration (RQ3) and different migration policies (RQ4). Since WarmSwap shows performance improvements only for functions with larger dependencies, we focus on the results for “lr_serving,” “cnn_serving,” and “rnn_serving” in this experiment.
Two migration policies of WarmSwap from Section 3.3 are evaluated:
WarmSwap Initiative Restore:
The page server copies all memory pages in parallel after the first “Page Fault.”
WarmSwap Lazy Restore:
The page server copies only the accessed pages that trigger a “Page Fault” at a time.
Two variants of WarmSwap are created as follows:
WarmSwap without Page Server:
Prepares a regular checkpoint image for the dependency, copies it to the user container during cold starts, and restores it.
WarmSwap without Lazy Migration:
Copies all memory pages through the page server inside the user container before the execution of user code.
Prototypes Cold Start (s) Warm Start (s)
lr_serving cnn_serving rnn_serving lr_serving cnn_serving rnn_serving
WarmSwap Bulk Restore 2.67 2.06 0.89 1.80 0.65 0.004
WarmSwap Lazy Restore 2.93 3.29 0.76 2.39 2.71 0.004
w/o Page Server 3.00 2.14 0.91 1.82 0.68 0.004
w/o Lazy Migration 3.09 2.13 0.89 1.81 0.65 0.004
Table 2: Comparison of cold-start and warm-start results with: WarmSwap bulk restore, WarmSwap Lazy Restore; WarmSwap without Page Server; and WarmSwap without Lazy Migration.
Table 2 shows the cold-start and warm-start times for WarmSwap and its variants. WarmSwap with bulk restore has the best latency for “lr_serving” and “cnn_serving”, while WarmSwap with lazy restore outperforms the others for “rnn_serving”. The cold-start performance difference between bulk restore and lazy restore is related to the number of page faults a function triggers. The overhead of lazy restore increases significantly as the number of page faults rises. This is because whenever the function triggers a page fault, it needs to pause and wait for the page server to install the specific memory pages.
For WarmSwap without Page Server, we need to copy a relatively large image to the user container and restore all memory pages before execution. The overhead of copying the image is related to the checkpoint sizes, which are listed in Table 3.
For WarmSwap without Lazy Migration, the cold-start time is slower than WarmSwap because it needs to transfer all the memory pages through the network. Note that for “lr _serving”, WarmSwap without Lazy Migration is even slower than WarmSwap without page server. This is because “lr _serving” has a smaller checkpoint image size than the other functions.
All three prototypes (not counting Lazy Copy) have similar warm start times, which demonstrates that those three migration methods do not impact the time for warm starts. For Lazy Copy, the performance of warm starts decreases because “lr_serving” and “cnn_serving” trigger extra page faults during the first warm start. The performance will become normal for the following warm starts.
Based on these experiments, we believe bulk restore is more suitable in practice since the provider usually does not know how many page faults the target program will trigger. Moreover, it has no impact on warm-start performance.
Functions Metadata Size (MB) Image Size (MB)
helloworld 0.88 8.1
json_dumps_load 1.2 16
pyase 0.88 8.3
chameleon 0.88 8.9
lr_serving 5.6 79
cnn_serving 15 190
rnn_serving 12 200
Table 3: The size of data that needs to be transferred from the worker node to the function container of WarmSwap and WarmSwap without Page Server.
4.5 Case Study: Dependency Image Sharing Across Serverless Functions
In this section, we conduct simulation experiments to show the benefit of WarmSwap on sharing dependency images across different serverless functions to answer RQ5. We modify the “rnn_serving” functions by adding new code (using NumPy and torch) to the original user handler to create 10 different serverless functions. Assuming these 10 serverless functions are less frequently invoked, there is at most one instance for each function. However, the provider would still wish to gain the benefits of warm start even on these functions, if the cost is reasonable, in order to reduce the cold-start overhead.
Figure 7 compares WarmSwap and a checkpoint image-based method, Prebaking [23]. We implement Prebaking following the description in their paper, by checkpointing the entire previous warm serverless container with CRIU [7] to a function image and restoring it during the cold start. Note that, unlike WarmSwap and Baseline, the cost of container creation is larger in the case of Prebaking. Therefore, Figure 7 also includes the overhead of container creation for both methods.
Figure 7: Comparison of the cold-start latency and required cache/memory of WarmSwap and Prebaking Image. The experiment is based on 10 different serverless functions that require the same dependency image (Python+numpy+torch).
Figure 7 shows the accumulated cold-start latency (the sum of average cold-start latency for the 10 functions), and the required cache in memory needed by the optimization. WarmSwap significantly outperforms Prebaking in the cold-start performance (4.8 times faster), since restoring a large container with CRIU is much slower than migrating only the dependency.
Furthermore, the provider only needs 260 MB memory to warm up the 10 different functions when using WarmSwap. In Prebaking, the provider needs to keep 10 checkpoint images (about 1.8 GB) in the cache for the 10 different functions. Hence, warming up the less frequently used functions with WarmSwap is cheaper than Prebaking.
Although function-specific tuning methods, like vHive [25] and Catalyzer [8], achieve faster container restore, the provider must create and keep all function images in cache to optimize the less frequently used serverless functions. Thus, WarmSwap provides a practical method for providers to optimize cold starts for all functions at a reasonable cost.
5 Related Work
5.1 Function-Specific Tuning
One straightforward way to perform function-specific tuning is by using the Checkpoint/Restore techniques like CRIU [7]. One can create a checkpoint image of a specific warm serverless function and restore that image instead of initializing it from scratch during a cold start. For example, Prebaking [23] utilized transparent checkpoint technology (CRIU) to locally checkpoint the serverless function. This approach involves restarting the process during cold starts.
Catalyzer, as presented by Du et al. [8], confronts the intricate task of swiftly restoring serverless sandboxes, containers, virtual machines, and more. This innovative approach adeptly reinstates a virtualization-based function instance from a structured checkpoint image. Moreover, Catalyzer advances this restoration process by enabling on-demand recovery of both user-level memory states and system states. Notably, Catalyzer introduces the concepts of “sandbox fork” and “fork boot” to effectively tackle the challenge of warm starts, particularly when one or more hot instances already exist.
Building on checkpoint images, researchers have proposed more advanced optimizations. The authors of Catalyzer extended their contributions through MITOSIS [26]. This work takes the notion of sandbox forking further by leveraging Remote Direct Memory Access (RDMA) for remote forking. This innovation enables rapid warm starts for multiple sandboxes and supports stateful serverless computing. This extension enhances the versatility and applicability of the sandbox forking concept, facilitating efficient warm start strategies and accommodating stateful serverless scenarios.
Building upon the foundational work of Catalyzer, another study[25] introduces a noteworthy approach termed “hot page record-and-prefetch technology” aimed at expediting sandbox restoration and serverless function execution. The underlying concept revolves around recognizing that not all data within the C/R image will be utilized following restoration.
Additionally, the SEUSS framework [6] presents a unique C/R-like solution. In contrast to other approaches, SEUSS focuses on the restoration of function unikernel snapshots, sidestepping resource-intensive initialization steps. An intriguing aspect of SEUSS is its incorporation of page-level sharing across the entire software stack required for function execution. SEUSS encompasses not only unikernel binaries but also runtime snapshots and user code snapshots.
There are more ways to conduct function-specific tuning to speed up cold starts by analyzing user function code. Fasslight [17] optimizes code loading latency by loading only indispensable code. To implement this, Fasslight constructs the function-level call graph, separates other code, and loads only indispensable code to reduce cold start overhead.
Function-specific tuning usually achieves extremely high-speed cold starts, as reported by MITOSIS [26] and vHive [25]. However, as we analyzed in Section 2, these methods are not suitable for all serverless functions in production, especially those that are less frequently used. Additionally, some methods are difficult to implement on the cloud provider side because they require the provider to record [25] or analyze user code [17] or containers.
5.2 Infrastructure Optimization
For infrastructure optimization, the easiest way is to use a better hardware infrastructure. Further, there are plenty of system designs that help faster loading and initializing the functions by virtualization and caching systems.
Firecracker [1] by AWS Lambda is a Virtual Machine Monitor that utilizes the Linux Kernel’s KVM virtualization infrastructure to offer lightweight virtual machines. Firecracker significantly reduces the boot time of serverless instances and enhances switching performance.
AWS further reports its container optimization, which has implemented sharing across serverless functions in its approach to loading containers on-demand [5]. This is achieved through transforming serverless containers into data chunks, which are then deduplicated [18, 27] and shared across functions. These data chunks are efficiently stored within a three-tiered cache system. The primary focus of this work is on optimizing storage and facilitating the swift movement of larger serverless containers, all aimed at accelerating cold-start times.
The inherent nature of serverless functions has opened avenues for sharing not only within the same function but also across different functions. This is particularly valuable for frequently utilized functions that may accumulate millions of instances. Developers often opt for identical containers, runtimes, or even packages. In a recent report, AWS revealed that a significant 80% of newly uploaded serverless containers lack new data chunks [5].
Building on this insight, researchers extend infrastructure optimization through shareable resources. Pagurus [14], for instance, explores container-level sharing. Recognizing that certain user functions extensively share software packages, this approach allows new functions to borrow containers from existing ones, thereby reducing cold-start overhead. However, it’s crucial to acknowledge that sharing entire containers could introduce additional security challenges. HotC [24] focuses on reusing container runtimes to minimize container initialization overhead. HotC maintains a live container runtime pool, analyzes user input or configuration files, and provides available runtimes for immediate reuse.
In this work, we propose a feasible solution to extend infrastructure optimization for reducing dependency loading overhead.
5.3 Dependency Loading Overhead
Several previous works have recognized the importance of software dependency loading overhead. The SOCK framework, as explored in Oakes et al.’s work [21], analyzed the impact of loading Python packages on overhead. Their findings revealed that incorporating numerous popular libraries during the import process adds approximately 100 milliseconds to the cold-start time. Consequently, they devised an innovative strategy involving the creation of a lean container and a three-tiered Python library caching system. This approach effectively accelerates cold-start times.
Similarly, the FlashCube framework [16] addresses the delays caused by initializing function runtimes. To tackle this issue, FlashCube introduces a novel container framework. Instead of constructing a container from scratch, FlashCube assembles it by combining a selection of pre-existing general container components, including namespaces, cgroups, and runtimes. This approach maintains a template runtime process for each programming language (such as Python, Java, and Node.js). Photons [9] also recognize language runtime and library redundancy. They co-locate multiple instances of the same function within the same runtime, allowing concurrent invocations to share dependencies and application states.
Note that existing solutions attempt to pre-warm dependencies at the container level, requiring providers to pre-warm containers or container components for serverless functions. In contrast, WarmSwap is a lightweight solution that can be easily deployed across multiple nodes while effectively addressing the dependency loading problem.
6 Conclusion
In conclusion, this paper presents WarmSwap, a provider-side solution for accelerating dependency loading in serverless cold starts in production serverless platforms. WarmSwap is a complementary work that extends infrastructure optimization in serverless environments by representing software dependencies as middleware. It is designed to benefit the many less frequently used serverless functions, which is often not practical in a function-specific approach. WarmSwap is primarily beneficial when the dependencies of the serverless functions are large and when they incur significant latency during initialization. Containers with short initialization times already perform well during cold start, but recent years have seen greater use of complex software dependencies like AI/ML. It is in this regime that WarmSwap is expected to be used.
By using WarmSwap, a provider can implement a single pool of dependency images across its many computer hosts. This relieves the user of additional work. Since dependency images represent public and commonly used middleware, they can be shared among a whole category of serverless functions. Unlike previous research, the provider only needs to pre-warm and migrate the dependency process itself instead of entire containers. We implemented a prototype of WarmSwap based on offline AWS Lambda containers, and the evaluation shows that WarmSwap reduces the dependency loading overhead and warm-up cost across many serverless functions.

### v3

@@
-WarmSwap: Sharing Dependencies for Accelerating Cold Starts in Serverless Functions
+Boston, MA, USA
 Abstract
-This work presents WarmSwap, a novel provider-side cold-start optimization for serverless computing.
+This work presents HotSwap, a novel provider-side cold-start optimization for serverless computing.
 This optimization reduces cold-start time when booting and loading dependencies at runtime inside a function container.
-Previous approaches to the optimization of cold starts tend to fall into two categories: optimizing the infrastructure of serverless computing to benefit all serverless functions; or function-specific tuning for individual serverless functions.
-In contrast, WarmSwap offers a broad middle ground, which optimizes entire categories of serverless functions.
-WarmSwap eliminates the need to initialize middleware or software dependencies when launching a new serverless container, by migrating a pre-initialized live dependency image to the new function instance.
-WarmSwap respects the provider’s cache constraints, as a single pre-warmed dependency image in the cache is shared among all serverless functions requiring that software dependency image.
-WarmSwap has been tested on seven representative functions from FunctionBench.
-The functions are chosen to compare with previous work.
-In those tests, WarmSwap accelerates cold-start executions for those serverless functions with large dependency requirements by a factor ranging from 1.2 to 2.2.
-1 Introduction
+Previous research has extensively focused on reducing cold-start latency for specific functions.
+However, little attention has been given to skewed production workloads.
+In such cases, cross-function optimization becomes essential.
+Without cross-function optimization, a cloud provider is left with two equally poor options: (i) Either the cloud provider gives up optimization for each function in the long tail (which is slow); or (ii) the cloud provider applies function-specific optimizations (e.g., cache function images) to every function in the long tail (which violates the vendor’s cache constraints).
+HotSwap demonstrates cross-function optimization using a novel pre-warming strategy.
+In this strategy, a pre-initialized live dependency image is migrated to the new function instance.
+At the same time, HotSwap respects the provider’s cache constraints, because a single pre-warmed dependency image in the cache can be shared among all serverless functions that require that image.
+HotSwap has been tested on seven representative functions from FunctionBench.
+In those tests, HotSwap accelerates dependency loading for those serverless functions with large dependency requirements by a factor ranging from 2.2 to 3.2.
+Simulation experiments using Azure traces indicate that HotSwap can save 88% of space, compared with a previous function-specific method, PreBaking, when sharing a dependency image among ten different functions.
+Index Terms:
+Cloud Computing, FaaS, Serverless Computing, Cold Start, Dependency Pre-Warming
+I Introduction
 Serverless computing is a flexible and popular cloud computing methodology that provides users with a simpler and cheaper way to deploy applications.
+A common example is Function-as-a-Service (FaaS), which enables event-driven execution of individual functions.
 It has quickly gained attention across various cloud providers, including AWS [1], Azure [2], Google[3], and Alibaba [4].
-For the user, serverless computing offers robust auto-scaling and an economical pay-as-you-go model, in contrast to reserving an instance for a fixed time.
-Users only need to submit their code as a serverless function in the cloud and later trigger the function’s execution based on a request or event.
-Finally, the user’s code is executed in instances (e.g., a virtual machine or a Docker container) provided by cloud providers.
 However, many of the financial benefits are undermined by the cold-start problem.
-In serverless computing, the cloud provider loads user function into the serverless instance only upon user invocation, forming the basis of the “pay-as-you-go” pricing model.
-Typically, the cloud retains the serverless instance in memory for a certain period (i.e., during a keep-alive time).
+In FaaS, the cloud provider loads user function into the serverless instance only upon user invocation.
+After an invocation, the cloud retains the serverless instance in memory for a certain period (i.e., during a keep-alive time).
 If an invocation occurs when no instance is alive in the system, the function experiences a cold start, which is significantly slower due to instance initialization.
-If the next invocation happens within the keep-alive time, the function experiences a warm start, during which the provider does not need to load and initialize a new instance.
-Cold starts can also occur during spikes in invocations when the system requires additional instances of a given serverless function.
-However, as reported by Azure [22], this affects only a small percentage of applications (less than 1%).
-Given the significance of the cold-start problem [10], recent years have witnessed a surge in related research efforts from both academia and industry [6, 8, 21, 23, 26, 20, 15, 24].
-To address the overhead of cold start, these researchers propose various system designs that leverage techniques such as caching, remote memory, and checkpoint/restore.
-Their ultimate goal is to achieve a faster cold start by accelerating or bypassing certain initialization phases.
-Optimizations for serverless cold start can be classified into function-specific tuning and infrastructure optimization based on their targets.
-Function-specific tuning focuses on individual functions, such as creating and caching pre-initialized images [23, 8, 6], generating function seeds [26], and recording and pre-fetching hot pages [25].
-Infrastructure optimization typically targets the entire system or shareable resources to benefit different functions.
-Examples include building a more efficient storage or caching system [5, 1], maintaining pre-initialized Zygote containers [21, 16], and sharing idle containers across different functions [14].
-Function-specific tuning typically results in very fast cold starts because it can manage every detail of the initialization phase for a specific function.
-However, implementing these optimizations for all serverless functions on the cloud can incur substantial costs for the cloud provider.
-In contrast, infrastructure optimizations have lower costs for large-scale cloud providers because they naturally benefit all serverless functions running on the cloud.
-Furthermore, infrastructure optimizations are easier to implement in real-world cloud environments because they only require involvement from the cloud providers.
-Figure 1: The relationship of invocation rate $\lambda$ versus number of expected cold starts with keep-alive time $T=15$ minutes and the function distribution coming from Azure [22].
+Otherwise, the function experiences a faster warm start with no initialization phase.
+Due to the significance and challenges of the cold start problem, it has been extensively studied by systems researchers [5, 6, 7].
+A naive solution to avoid the cold start is to always keep the function instance alive, ensuring that cold starts never occur.
+Researchers have further developed approaches to intelligently adjust the keep-alive duration based on invocation patterns [8, 9] or by predicting invocation arrivals and pre-loading functions in advance [10, 8].
+However, most user request arrivals are hard to predict [11, 12].
+Keeping idle functions alive wastes resources, prompting research into accelerating cold starts instead.
+A common approach is to pre-warm functions using checkpoint images or caches [13, 14, 15, 16], which bypass parts of the initialization process.
+While effective, these methods increase memory usage, straining limited cache resources.
+Recent work refines this by pre-fetching hot pages [17] and generating function seeds [18], but further optimization is needed to reduce overhead and improve performance.
+Experiments from a recent state-of-the-art work show that these methods improve the cold-start tail latency by 89% and require only 29MB of memory per machine to pre-warm a serverless function [18].
+These methods optimize cold starts by analyzing function-specific initialization bottlenecks.
+However, they are not lightweight, often requiring deep analysis or specialized OS and hardware support [18], such as RDMA.
+Nevertheless, according to Azure Function traces [11], 50% of functions experience fewer than $1.4$ cold starts per day, assuming a keep-alive time of 15 minutes.
+Therefore, maintaining pre-initialized images or recording hot pages in RAM for every serverless function in a workload like Azure [11] is costly and unnecessary.
+Implementing function-specific optimizations for all serverless functions is impractical in production due to skewed workloads, where many functions experience few cold starts (see Figure 1).
+For functions with invocation rates far from $1/T$, the cost of applying prior methods is not economical, where $T$ is the keep-alive time.
+A more practical approach, albeit with other limitations, is cross-function pre-warming, where shared resources are preserved across functions [19, 20, 21].
+For instance, SOCK [19] and FlashCube [21] reduce startup latency respectively by retaining Zygote containers or by retaining shared components on worker nodes.
+However, maintaining or migrating such pre-initialized containers remains costly at scale.
+Figure 1: The relationship between invocation rate $\lambda$ and expected cold starts with keep-alive time $T=15$ minutes, using function distribution from Azure [11].
 Each serverless function was assigned to a bucket $[x,x+0.001]$, where the $\lambda$ of the serverless function was between $x$ and $x+0.001$ invocations per minute.
-The y-value for a bucket is the number of functions that fall into the given bucket.
-The y-values for each bucket were normalized so that the sum of the values for the buckets adds to one.
-In practice, production cloud providers tend to emphasize more infrastructure optimization, due to the skewed workload in the real world, as illustrated in Figure 1.
-Function-specific tuning is suitable for functions with invocation rates around $1/T$, where a significant number of expected cold starts occur to each function.
-However, according to an Azure Function dataset [22], 50% of functions experience fewer than $1.4$ cold starts per day, assuming a keep-alive time of 15 minutes.
-Therefore, it is costly and unnecessary to maintain pre-initialized images or record hot pages in RAM for all serverless functions.
-Although some infrastructure optimization for serverless container caching and launching have been successfully implemented in production [1, 5], there remains a lack of infrastructure optimization for loading software dependencies.
-For instance, AWS Lambda can launch cold starts for function containers up to 10 GB in size within 50 ms [5], but it still initializes software dependencies from scratch (see experiments in Section 2.3).
-The time required to load software dependencies within the container can sometimes exceed the execution time [21].
-Recently, this initialization time has been exacerbated by the development of AI and large language models, as AI-related functions often require larger and more complex dependencies.
-The existing work targets reducing dependency-loading latency, but deploying those solutions in real-world serverless systems still presents challenges.
-SOCK [21] and FlashCube [16] have proposed new container frameworks that aim to bypass container launch and dependency-loading phases by preserving Zygote containers and container components on serverless worker nodes.
-However, preserving pre-initialized containers or their components on every serverless worker node, or remotely migrating the entire container, is costly.
-It also restricts user flexibility to use their own custom containers, a feature commonly supported by large serverless providers like AWS [3].
-Finally, this work is complementary to previous work, and can effectively coexist with previous approaches.
-The function-specific tuning like catalyzer [8] or vHive [25] optimize the frequently used serverless functions (i.e.
-around $1/T$ in Figure 1).
-This work optimizes the “long tail” of the distribution of buckets for serverless functions (see caption of Figure 1), as documented by Azure [22].
-And finally, infrastructure optimizations benefit [5, 1] all serverless functions.
-The WarmSwap Approach:
-In this paper, we present WarmSwap, targeting the dependency loading overhead during cold starts.
-In WarmSwap, the provider maintains a pool of live dependency images in RAM.
-A live dependency image is an image stored in memory that represents a checkpoint of the software dependencies.
-Typically, these images contain popular middleware combinations, such as Python with NumPy, Python with Torch, or Python with NumPy and Torch.
-WarmSwap is designed to benefit serverless providers by also optimizing the massive number of less frequently invoked functions (the “long tail” of the distribution), which is usually not practical in a function-specific approach.
-WarmSwap brings several benefits.
-First, providers only need to maintain a limited number of live dependency images in memory, which can be shared by a wide variety of serverless functions that utilize that middleware.
-The pre-warming cost for the provider depends solely on the number of live dependency images, not on the number of serverless functions employed by multiple users.
-However, the benefit of pre-initialization is shared among all serverless functions using the middleware.
-Second, WarmSwap respects isolation for serverless functions.
-This is because the provider primarily transfers pages from public dependency images, which do not include user source code.
-By using WarmSwap, the provider can accelerate the cold start by “pre-warming” a popular dependency pool in either the local worker nodes or a remote central node.
-When creating new serverless instances for cold starts, WarmSwap can first check whether a necessary software dependency already exists in the dependency image pool.
-If so, WarmSwap conducts a live migration of the required dependency image from the pool to new serverless containers.
-This saves time otherwise spent on booting runtimes and loading software dependencies.
-Furthermore, the WarmSwap approach is practical because it focuses on widely used, public software dependencies.
-A study [21] reveals that 36% of “import” in 876,000 Python projects are from just 20 packages.
-Therefore, providers can concentrate on maintaining a relatively small number of dependency images across numerous serverless functions used by all users.
-An experimental implementation of WarmSwap was evaluated by running offline AWS Lambda containers on AWS EC2 nodes.
-In our tests, loading libraries such as “sklearn+pandas”, “numpy+keras”, and “numpy+torch” through WarmSwap was up to 2.2x, 3.2x, and 2.5x faster, respectively, compared to traditional AWS Lambda startup processes.
-Furthermore, WarmSwap can save 85.4% of memory space compared to the function-specific approach [23] when warming up 10 different serverless functions sharing the same dependency.
-Therefore, for functions sharing a common dependency, WarmSwap represents a new type of infrastructure optimization using pre-warmed dependency images, offering several advantages over previous approaches in the literature.
+The y-value for a bucket is the normalized number of functions in that bucket.
+Despite extensive research on the cold start problem, key challenges remain and hinder large-scale deployment in production.
+In particular, there is a lack of lightweight, cost-efficient, cross-function optimizations that target the majority of infrequently invoked functions (see the caption of Figure 1).
+This paper aims to enable cross-function optimizations in the dependency-loading phase, a major bottleneck in current production cloud environments (see Section II-C), while keeping overhead minimal.
+Thus, we need only apply cross-function optimization to the frequently used dependencies instead of entire images, thereby improving scalability.
+The HotSwap Approach:
+In this paper, we present HotSwap, a cross-function solution targeting the dependency-loading overhead during cold starts.
+HotSwap enables the provider to maintain a pool of live dependency images in RAM.
+A live dependency image is a memory-resident checkpoint representing software dependencies.
+Typically, these images include popular software dependency combinations, such as NumPy and Torch.
+HotSwap is designed to benefit serverless providers by also optimizing the massive number of less frequently invoked functions (the “long tail” of the distribution), which is usually not practical in a function-specific approach.
+HotSwap brings an important benefit.
+Providers only need to maintain a limited number of live dependency images in memory, which can be shared across a wide range of serverless functions using that dependency.
+Therefore, for functions sharing a common dependency, HotSwap represents a new type of cross-function optimization using pre-warmed dependency images, offering several advantages over previous approaches in the literature.
+Supporting this, a study [19] analyzed 876,000 Python projects and found that just 20 packages accounted for 36% of all ‘import’ statements, indicating that most functions rely on a small set of common dependencies.
+A prototype of HotSwap was evaluated by running offline AWS Lambda containers on AWS EC2 nodes.
+The implementation of HotSwap, unlike some other approaches, is entirely transparent, without modification to the operating system kernel or the source code of the serverless function.
+In our tests, loading libraries such as “sklearn + pandas”, “NumPy + keras”, and “NumPy + torch” through HotSwap was up to 2.2x, 3.2x, and 2.5x faster, respectively, compared to traditional AWS Lambda startup processes.
+Furthermore, experiments in Section IV-E using traces from Azure [11] show that HotSwap can save 88.4% of memory space compared to a function-specific approach, PreBaking [13, 14].
+This efficiency is achieved when warming up 10 different serverless functions that share the same dependency.
 We summarize our contributions as follows.
   1. 1.
-    This paper proposes WarmSwap, a new infrastructure optimization targeting the loading of a software dependency to accelerate serverless cold starts, benefiting the many less frequently invoked serverless functions as well.
+    This paper proposes HotSwap, a novel cold-start optimization targeting skewed, long-tail serverless workloads, which are common in commercial clouds like Azure.
   2. 2.
-    WarmSwap is a cost-effective solution for dependency-loading, as the provider stores middleware only once in a separate dependency image, independent of the entire container.
-A single pre-warmed dependency image can be shared between different serverless functions and worker nodes.
+    HotSwap introduces a new solution to the dependency-loading issue, where the provider stores middleware once in a separate dependency image, independent of the entire container.
+A single pre-warmed dependency image can be shared across different function instances.
   3. 3.
-    This work is targeted to benefit the real-world distributions of serverless function invocations with their “long tail”, as documented by an Azure study (see Figure 1).
-The rest of this paper is organized as: Section 2 presents the background and motivation of our paper; Section 3 presents the design and implementation details of WarmSwap; In Section  4, we evaluate the performance of WarmSwap; We then discuss related work in Section 5, and conclude in Section 6.
-2 Background and Motivation
-This section introduces the background for serverless computing and the issue of cold starts in Section 2.1.
-Section 2.2 motivates the need for WarmSwap by analyzing the relationship between cold starts and function invocation patterns in Azure.
-Section 2.3 then discusses an existing performance bottleneck in a commercial cloud platform such as AWS Lambda.
-2.1 Cold Start in Serverless Computing
+    HotSwap was evaluated for its ability to reduce provider-side costs (CPU usage, cache, RAM) during cold starts using a real-world workload from Azure [11], rather than a synthetic workload.
+The rest of this paper is organized as: Section II presents the background and motivation of our paper; Section III presents the design and implementation details of HotSwap; In Section  IV, we evaluate the performance of HotSwap.
+We then discuss related work in Section V, and conclude in Section VII.
+Figure 2: The process of serverless cold starts in the current production cloud
+II Background and Motivation
+This section introduces: the background for serverless computing and cold starts (Section II-A); the need for HotSwap (Section II-B); and a performance bottleneck in a commercial cloud platform such as AWS Lambda (Section II-C).
+II-A Cold Start in Serverless Computing
 This section introduces the background of container-based serverless computing from a cold-start perspective. Figure 2 shows the general process of such a cold start.
-Figure 2: The process of serverless cold starts in the current production cloud: Different optimization methods may involve one or more steps in the process.
-In Figure 2, while the specifics may differ across platforms, most serverless architectures typically consist of two core components: a Worker Manager and a Resource Orchestration unit.
-In response to a function invocation from the front end, a request for capacity is sent to the Worker Manager, which acts as a stateful load balancer.
-The Worker Manager continuously monitors the capacity of each unique function on the platform.
+In Figure 2, while the specifics may differ across platforms, most serverless architectures consist of two core components: a Worker Manager and a Resource Orchestration unit.
+The Worker Manager responds to a request for capacity from the serverless front end.
+The Worker Manager acts as a stateful load balancer by continuously monitoring the capacity of each unique function on the platform.
 If sufficient capacity exists (e.g., idle function instances are available), the Worker Manager directs the front end to route the request to an idle instance.
-Conversely, if resources are unavailable (triggering a cold start), it instructs Resource Orchestration to provision a new function instance with adequate CPU and RAM.
+Conversely, if resources are unavailable (triggering a cold start), it instructs the Resource Orchestration unit to provision a new function instance with adequate resources.
 Recall also that the user has uploaded the container function to the provider’s Container Registry.
@@
 Then, the serverless runtime starts a specific function runtime (e.g., Python), and the program loads the necessary software dependencies (e.g., NumPy) and the user-specific handler code.
-At this point, the new function instance is ready to handle user requests.
-To minimize the cold-start overhead, one needs to skip or expedite one or more steps in the cold-start optimization.
-When designing a function-specific tuning for the cold start, one can consider all the steps specific to a serverless function.
-Previous works have designed end-to-end, function-specific tuning for targeted serverless functions and achieved substantial reductions in cold-start latency.
-However, when performing infrastructure optimization for the entire system, one can only optimize steps that are identical or similar across all serverless functions, such as container storage and loading.
-In this paper, we argue that runtime booting and dependency-loading qualify for infrastructure optimization because they are common processes that can benefit multiple serverless functions.
-Thus, we propose WarmSwap as a novel infrastructure optimization aimed at reducing the time required for runtime booting and dependency-loading.
-2.2 Impact of Invocation Patterns on Cold Starts
-In this section, we aim to address two questions:
-  1. 1)
-    How frequently does a cold start occur for a serverless function?
-  2. 2)
-    How does a skewed workload impact the optimization choices of the cloud provider?
+The handler code is then ready to handle a new function instance.
+To minimize the cold-start overhead, one needs to skip or expedite one or more steps.
+When designing a solution for a specific serverless function, all function-specific steps can be considered.
+Previous work has designed end-to-end, function-specific optimization for targeted serverless functions and achieved substantial reductions in cold-start latency.
+However, when performing cross-function optimization for the entire system, one can only optimize steps that are identical or similar across all serverless functions, such as container storage and loading.
+In this paper, we explore how runtime booting and dependency loading can benefit from cross-function optimization, as they are common processes shared by multiple serverless functions.
+Thus, we propose HotSwap as a novel cross-function optimization aimed at reducing the time required for runtime booting and dependency loading.
+II-B Impact of Invocation Patterns on Cold Starts
+In this section, we aim to address two questions: 1) How frequently does a cold start occur for a serverless function?
+and 2) How does a skewed workload impact the optimization choices of the cloud provider?
 To simplify the discussion, we focus on situations where a cold start occurs due to the absence of any active instances, as this scenario accounts for more than $99\%$ of cold starts in production systems [11].
@@
 Let $\lambda$ denote the invocation rate (calls per minute), and $T$ denote the keep-alive time.
-The probability of no invocation occurring during the keep-alive time follows an exponential distribution:
+The probability of no invocation occurring during the keep-alive time follows:
 $P_{no\_inv}=e^{-\lambda T}$ (1)
-Here, $P_{no\_inv}$ represents the probability that no invocation happens during the keep-alive timeout, leading to a potential cold start.
+$P_{no\_inv}$ represents the probability that no invocation happens during the keep-alive timeout, leading to a potential cold start.
 The number of expected cold starts that occur within $D$ minutes can be expressed as:
@@
 which has a maximum when $dE/d\lambda=0,\lambda=1/T$.
-Figure 1 shows the relation of $E_{cs}$ versus $\lambda$ when $T=10\;min$ and $D=1440\;min$ (a day).
+Figure 1 shows the relation of $E_{cs}$ versus $\lambda$ when $T=15\;min$ and $D=1440\;min$ (a day).
 Only functions with invocation rates around $1/T$ have a large number of expected cold starts, functions that are less or more frequently invoked have fewer cold starts.
@@
 The different keep-alive time changes the position of the maximum point, but the shapes are similar.
-To answer question two, we draw the density histogram of the actual function distribution in the same Figure 1.
+To answer question 2, we draw the density histogram of the actual function distribution in the same Figure 1.
 For function-specific tuning, assuming the cost of optimizing a function is $c$ and the benefits of an optimized cold start are $w$.
@@
 However, from the histogram, the serverless workload in production can be extremely skewed.
-Over $50\%$ of functions have very a low invocation rate (less than 0.001 calls per minute) and $E_{cs}$, which experience less than $1.4$ times of cold starts.
-Thus, those functions are hard to qualify equation 3, which means function-specific tuning is not economic for them.
-For those functions whose invocation rate cannot qualify equation 3, only infrastructure optimization can help to reduce their cold-start overhead.
-One successful optimization example from the provider is AWS Lambda, which published its optimization of function creation and container loading in a recent paper [5].
-The design of AWS lambda assumes that serverless instances are highly cacheable, in which most scaling is driven by a smaller number of images, and the images are based on common base layers.
-With a three-tier caching system and container deduplication, AWS can cold start a 10 GB serverless container within 50 ms.
-Thus, the loading and booting of the container are lightweight in current commercial serverless systems.
-In this paper, we hope WarmSwap can further extend the infrastructure optimization for the benefit of the many serverless functions that can not apply function-specific tuning.
-2.3 Performance Analysis of the Cold Start for Current AWS Lambda Functions
+More than $50\%$ of the functions have a very low invocation rate (less than 0.001 calls per minute) and $E_{cs}$, which experience less than $1.4$ times of cold starts per day.
+Thus, those functions do not satisfy Equation 3, which means function-specific tuning is not economical for them.
+For functions whose invocation rates do not meet the tuning criterion in Equation 3, only cross-function optimizations are feasible.
+One successful example is AWS Lambda [22], which employs a three-tier caching system and container deduplication to load a 10 GB serverless container in just 50 ms.
+Such optimizations are applicable to all serverless functions, regardless of invocation frequency.
+In this paper, we aim for HotSwap to further enhance cross-function optimization, benefiting serverless functions where function-specific tuning is not economically viable.
+II-C Performance Analysis of the Cold Start for Current AWS Lambda Functions
+Figure 3: Cold-start latency breakdown for serverless functions in a traditional AWS Lambda function (not HotSwap).
 This section analyzes the performance of cold-start overhead in a production cloud on AWS [23, 22].
 The goal is to analyze which components of the current AWS Lambda architecture still have room for optimization.
-None of the experiments in this section refer to WarmSwap.
-For analysis, we built serverless function containers to create AWS Lambda functions, and then invoke them by a script running on an AWS EC2 node.
+For analysis, we build serverless function containers to create AWS Lambda functions, and then invoke them by a script.
 All Lambda functions are configured with a 10 GB memory limitation to maximize the CPU resources.
-To collect cold-start times, the script needs to wait until the end of keep-alive time between invocations of the same function.
+To collect cold-start times, we delete and recreate lambda functions after each run.
 There are several timers in the analysis script and the function containers to track the time of each cold-start process.
 The entire time for the cold start is calculated from the script that sends the request to the serverless function and receives the result.
-Note that the caching system of AWS Lambda would bring a bias to a naive experimental design.
+Note that the caching system of AWS Lambda would cause bias in a naive experimental design.
 For example, if the user uploads a new function container, the first cold start will be extremely slow as compared to the following ones.
 Moreover, if the functions are frequently invoked, AWS upgrades the function to a different cache that will improve its cold starts.
-(See “AWS three-tier caching” in [5].) The improved caching scheme continues to be used even when the function is deleted and then recreated.
-In order to achieve reproducible numerical results, we upload a container and then invoke the serverless function many times, until the cold-start results are stable.
-Serverless Functions Container Size (MB) Major Dependencies
-helloworld 180.00 None
-json_dumps_load 180.00 urllib, json
-pyase 180.11 pyaes, string
-chameleon 180.42 chameleon
-lr_serving 379.52 sklearn, pandas
-cnn_serving 1386.73 numpy, keras
-rnn_serving 5602.46 numpy, torch
-Table 1: Serverless functions adopted from FunctionBench [12]
-Table 1 presents serverless functions from a representative serverless benchmark suite called FunctionBench [12, 13].
-We have chosen FunctionBench, for ease of comparison with previous work [25].
+(See “AWS three-tier caching” in [22].) The improved caching scheme of AWS Lambda continues to be used even when the function is deleted and then recreated.
+In order to achieve reproducible numbers with lower variance in our own experiments on AWS Lambda, we upload a container and then create and delete a Lambda function many times, until the cold-start results are stable.
+TABLE I: Serverless functions adopted from FunctionBench [24].
+Serverless Functions Container Size (MB) Major Dependencies Function Purpose
+helloworld 180.00 None Minimal function
+json_dumps_load 180.00 urllib, json Json Serialization
+pyase 180.11 pyaes, string Text encryption
+chameleon 180.42 chameleon HTML rendering
+lr_serving 379.52 sklearn, pandas Logistic regressions
+cnn_serving 1386.73 NumPy, keras Image classification
+rnn_serving 5602.46 NumPy, torch Sequence generation
+Table I presents serverless functions from a representative serverless benchmark suite called FunctionBench [24, 25].
+We have chosen FunctionBench for ease of comparison with previous work [17].
 All these functions are written in Python and they use some representative Python packages like “NumPy”, “torch”, “keras”, and “sklearn”.
-Table 1 lists both packages and the size of specific function container images.
-Installing large packages also significantly increases the image size of function containers.
-Section 4.1 describes more details of these functions.
-Figure 3: Cold-start latency breakdown for serverless functions in a traditional AWS Lambda function (not WarmSwap).
-Figure 3 shows the breakdown of intermediate times for cold starts in this analysis. The entire cold-start time can be broken down into to following parts:
+Table I lists both packages and the size of specific function container images.
+Installing large packages also significantly increases the image size of the function containers.
+Figure 3 shows the breakdown of intermediate times for cold starts in this analysis. The entire cold-start time can be broken down into the following parts:
   • Network: The process of network communications between invoker and user functions.
@@
 From Figure 3, the dependency initialization dominates the cold start of AWS Lambda functions.
-In “lr_serving”, “cnn_serving”, and “rnn_serving”, the dependency initialization occupies $43\%$, $74\%$ and $68\%$ of the cold-start overhead, respectively.
-The time for the dependency initialization component is even higher than that for the execution component for the three AWS Lambda functions.
-The dependency initialization component is a major problem for serverless functions in production.
-This problem occurs frequently because the packages they need are so common.
-Note that the time for the container component is almost constant (about $0.5~{}s$), even for containers that have different image sizes.
-This shows that AWS has already optimized the time for the container component, by improving the AWS caching and file systems [5, 1].
-In summary, the time of the dependency initialization component often dominates the time for the other components.
-Therefore, improving the dependency initialization component overhead is important both to serverless providers and to users of the current production cloud.
-This motivates a novel method focusing on the dependency initialization overhead, as seen in Section 3.
+In “lr_serving”, “cnn_serving”, and “rnn_serving”, the dependency initialization occupies $43\%$, $74\%$, and $68\%$ of the cold-start overhead, respectively.
+Dependency initialization takes more time than execution for all three AWS Lambda functions.
+It remains a major bottleneck for serverless functions in production, especially since many rely on widely used packages.
+Note that the time for container creation is almost constant (about $0.5~{}s$), even for containers that have different image sizes.
+This shows that AWS has already optimized the time for the container creating, by improving the AWS caching and file systems [22, 23].
+In summary, the time for dependency initialization often dominates the cold-start time in AWS Lambda.
+Therefore, improving the dependency initialization overhead is important both to serverless providers and to users of the current production cloud.
+This motivates a novel method focusing on the dependency initialization overhead, as seen in Section III.
 (a) System Architecture
@@
 (c) Runtime Phase
-Figure 4: System architecture and workflow: (a) Architectureal overview of WarmSwap and its dependency pool.
-(b) In the setup phase, the user registers a serverless function, and WarmSwap generates the corresponding image in the dependency pool.
+Figure 4: System architecture and workflow: (a) Architectural overview of HotSwap and its dependency pool.
+(b) In the setup phase, the user registers a serverless function, and HotSwap generates the corresponding image in the dependency pool.
 (c) In the runtime phase, pages are copied from the image between provider worker nodes and user function containers.
-3 WarmSwap Design
-This section introduces WarmSwap, a prototype system for enabling “live dependency image” on serverless platforms.
-WarmSwap is motivated by the goal of improving the cold start by skipping the dependency loading phase.
-WarmSwap enables a serverless function container to seamlessly migrate the dependency image of its pre-initialized software dependencies from a provider’s worker node.
-For example, assume there is a serverless function based on Python and numpy.
-The function can use pre-initialized memory pages of Python and numpy through migration.
-This eliminates the otherwise large runtime overhead of first initializing Python and numpy.
-WarmSwap leverages the power of a shared dependency pool, enabling significant performance gains for providers.
-Even though individual functions differ, multiple users whose functions have a common dependency can benefit from the same pre-initialized dependency image from the pool.
-The cloud provider can then support such a dependency pool for all serverless users.
-In general, the pool contains several pre-initialized dependency images for popular runtimes and packages (e.g., Python + numpy, Python + torch, etc.).
-We first provide an overview of the system’s architecture in Section 3.1, highlighting the key components and their interactions.
-Next, we detail the implementation of each WarmSwap component, including the generation of the pool of dependency images in Section 3.2 and the design choices of migration in Section 3.3.
-Finally, Section 3.4 describes the implementation.
-3.1 Design Overview
-WarmSwap is designed to reduce dependency loading overhead for serverless functions.
+III HotSwap Design
+This section introduces HotSwap, a prototype system for cross-function cold-start optimization.
+HotSwap leverages the power of a shared dependency pool, enabling significant performance gains for providers.
+A single global shared dependency pool for all users is implemented for cross-function optimization.
+Hence, a single pre-initialized dependency image from the pool can benefit multiple users whose functions share a single common dependency.
+In general, the pool contains several pre-initialized dependency images for popular runtimes and packages (e.g., NumPy, torch, etc.).
+Thus, the cloud provider supports such a dependency pool for all users.
+HotSwap enables a serverless function container to seamlessly migrate the dependency image of its pre-initialized software dependencies from a provider’s worker node.
+For example, assume there is a serverless function based on NumPy.
+The function can use pre-initialized memory pages of NumPy through migration.
+This eliminates the otherwise large runtime overhead of first initializing NumPy.
+III-A Motivation for the Design of HotSwap
+HotSwap gains its performance advantages by exploiting three characteristics of the workloads seen by public cloud providers in their distribution of containerized serverless functions.
+Those three characteristics are:
+Commonality:
+Although there are numerous serverless software dependencies, and applications can use different combinations of dependencies, many applications are based on a small number of popular packages [19] (e.g., NumPy, torch).
+Sparsity:
+Most applications do not use all functions from a package during execution [17]. Thus, copying all memory pages to the function container beforehand is unnecessary.
+Lightweight:
+Keeping runtime and software dependencies in the memory is much more lightweight than keeping the whole container like [19]. Thus, the provider can afford to maintain a pool of dependency images.
+III-B Design Overview
+HotSwap is designed to reduce dependency loading overhead for serverless functions.
 The key idea is to migrate pre-initialized dependencies instead of starting them from scratch.
@@
 The system design is based on the characteristics of such dependencies.
-A container for a serverless function usually contains both user-specific code and a middleware platform with its own package dependencies.
-We refer to this middleware as a serverless software dependency.
-These containers generally share three characteristics, described below, which determine the design and performance of the WarmSwap workflow.
-Commonality:
-Although there are numerous serverless software dependencies, and applications can use different combinations of dependencies, many applications are based on a small number of popular packages (e.g., NumPy, sklearn, torch).
-Providing migration services for those popular packages can benefit complex serverless functions with large dependencies.
-Sparsity:
-Most applications do not need all functions from a package, which means they only need access to small parts of the dependency memory pages.
-In this situation, it is unnecessary to load all the memory pages before the execution of user functions.
-Lightweight:
-Keeping runtime and software dependencies in the memory is much more lightweight than keeping the whole container.
-Thus, the provider can afford to maintain a pool of dependency images on their work node.
-Figure 4a describes the components of WarmSwap on the provider work node and function container.
-On the provider worker node, a Dependency Manager orchestrates the management of in-memory or "live" dependency images, which are images stored in memory that represent a checkpoint of pre-initialized software dependencies (e.g., Python + numpy).
+A container for a serverless function usually contains both user-specific code and its software dependencies.
+Figure 4a describes the components of HotSwap on the provider worker node and function container.
+On the provider worker node, a Dependency Manager orchestrates the management of in-memory dependency images, which are images stored in memory that represent a checkpoint of pre-initialized software dependencies.
+The Dependency Manager maintains metadata about existing live dependency images and manages the memory pages associated with them.
 It is the central hub for communication with user function containers, facilitating migration services.
-The Dependency Manager maintains metadata about existing live dependency images and manages the memory pages associated with them.
 When a migration request is initiated by a serverless container, the Dependency Manager accepts the request, transfers the essential metadata required for process restoration, and provides page transfer services to the serverless function container.
 In the container for the user function, a Migration Client orchestrates the migration of memory pages on the user side.
-Upon startup of the container, it initiates a migration request for the Dependency Manager.
+Upon startup of the container, it initiates a migration request to the Dependency Manager.
 Upon receiving confirmation, it starts the Dependency Loader to orchestrate the restoration of the runtime and the loading of dependency images, guided by the provided metadata.
 Rather than loading all memory pages immediately, the Dependency Loader restores the memory pages of the dependency by lazy migration.
-The Dependency Loader restores the dependency without loading the actual memory pages; instead, it initiates a Page Server to handle this task later.
+In lazy migration, instead of the actual memory pages, the Dependency Loader initializes a Page Server to handle this task later.
 When the user functions execute and access memory pages of the dependency, it triggers page faults, and the Page Server handles them.
-This Page Server communicates with the dependency pool to efficiently load specific pages whenever required by the process.
-3.2 Generation of Dependency Images
-Figure 4b shows the setup phase of WarmSwap, which dumps a serverless software dependency into a dependency image in the dependency pool.
-The serverless software dependency process (refer to the target process in the following paragraphs) in this example is a Python script importing numpy and then executing the user handler.
+This Page Server communicates with the dependency pool to efficiently load memory pages.
+III-C Generation of Dependency Images
+Figure 4b shows the setup phase of HotSwap, which dumps a serverless software dependency into a dependency image in the dependency pool.
+The serverless software dependency process in this example is a Python script importing NumPy and then executing the user handler.
 We initiate the dumping of the process after NumPy has been loaded but before the import and execution of the user code (the user handler in Figure 4a).
-This dump timing ensures that the migrated process can support different user handlers and execute them immediately after migration.
-The target process is divided into two distinct components during the dumping process: process metadata, preserved within Dependency Metadata; and hot memory pages, residing securely within the Dependency pool.
+In general, other shared dependency images might also be used besides NumPy.
+The corresponding user handler will then execute immediately after that image has been migrated.
+The dependency process is divided into two distinct components during the dumping process: (i) process metadata, preserved within Dependency Metadata; and (ii) hot memory pages, residing securely within the Dependency Pool.
 The process metadata encompasses essential information for restoring the process framework, including file descriptors, memory segment sizes, and other relevant configuration details.
-The dependency pool preserves the full memory pages of the target process within its memory domain.
-To enable efficient page transfer, the Dependency Manager creates a dedicated page server process on the source for each target process.
-The source page server uses the network to communicate with page servers on the destination that need the target process.
-To further accelerate the process generation and optimize resource utilization, we employ checkpoint images on disk.
-When dumping target processes, we create checkpoint images in order on disk to further optimize the generation process.
-With images on disk available, we do not need to execute the Python script from the beginning.
-Instead, we can directly restore the checkpoint and then create live dependency images.
-This is useful for those large and less popular target processes because the provider can save costs by putting them on disk instead of memory.
-WarmSwap requires only a modest pool of dependency images, because of the commonality and lightweight characteristics of serverless software dependencies.
-Thus, cloud providers only undertake limited costs associated with dumping and maintaining dependent images.
-3.3 Migration of Dependency Images
-Figure 4c describes the runtime details of initializing a WarmSwap container.
-Due to the sparsity of the memory pages required by a dependency image, we leverage a strategy known as lazy migration.
-This approach dynamically transfers pages only as needed, significantly reducing overhead and enhancing performance.
+The dependency pool preserves the full memory pages of the dependency process within its memory domain.
+The process metadata, combined with its memory pages, is referred to as a dependency image in this work.
+To enable efficient page transfer, the Dependency Manager creates a source-side page server for each dependency image, which communicates over the network with destination page servers.
+HotSwap requires only a modest pool of dependency images, because of the commonality and lightweight characteristics of serverless software dependencies.
+Thus, cloud providers only undertake limited costs associated with dumping and maintaining dependency images.
+III-D Migration of Dependency Images
 The migration client orchestrates the migration process, which has three distinct steps.
-The first step of migration is to transport process metadata to the destination (i.e., to the serverless function container).
+The first step of migration is to transport the process metadata to the destination (i.e., to the serverless function container).
 After receiving the migration request from the migration client, the dependency manager looks up the dependency metadata for the requested dependency image.
@@
 The second step of migration is to start a page server, which handles the “page faults” of the dependency to be loaded by using the “userfaultfd” feature of Linux.
-The page server accesses the dependency pool through the network, to load those pages that actually are needed by the serverless function.
+The page server accesses the dependency pool through the network, to load those pages that are actually needed by the serverless function.
 In the third step, the dependency loader then restores the serverless software dependencies based on the process metadata.
@@
 After migration, the new process continues to import and execute the user handler.
-However, the dependency loader does not load all of the memory pages in the dependency image.
-Instead, it starts a page server that connects to a specific dependency image inside the dependency pool on the worker node.
-Thus, when the process accesses specific memory pages, it triggers a “Page Fault”, which the page server handles using “userfaultfd”.
-For handling the “Page Fault”, WarmSwap could implement either of two possible design choices:
+For handling the “page fault”, HotSwap could implement either of two possible policies:
 Lazy restore: When handling the “page fault”, only the pages required by the user functions are transferred.
@@
 Bulk restore: When handling the first “page fault”, the Page Server first transfers the required pages.
-And then, in the background, it loads the remaining pages in parallel until another page fault occurs.
+And then, in the background, it loads the remaining pages in parallel.
 In this case, the Page Server needs to access the dependency pool only once.
@@
 Although transferring all pages increases communication overhead, this should not impact the user functions because the Page Server can work in parallel with the user functions.
-WarmSwap chooses bulk restore because it generally benefits all functions without impacting the warm starts (see Section 4.4).
-In contrast, lazy restore only works well in fewer page fault cases.
-In a third possibility, one might record the subset of pages that the serverless functions need and only transfer them.
-For example, previous work [25] records and pre-fetch hot pages for each target serverless function.
-However, recall that this method involves function-specific tuning, which does not fulfill the requirements of WarmSwap.
-3.4 Implementation
-We implement a prototype of WarmSwap with AWS Lambda containers and CRIU [7] (Checkpoint and Restore In Userspace).
-WarmSwap uses AWS Lambda container templates (Version Python 3.9) to build serverless functions and collect performance results.
+HotSwap chooses bulk restore because it generally benefits all functions without impacting the warm starts (see Section IV-D).
+In contrast, lazy restore works well only in more specialized page fault cases.
+(a) Cold Start
+(b) Warm Start
+Figure 5: Comparison of HotSwap versus Baseline.
+The software dependencies for each function are specified in Table I of Section II-C.
+Note that lr_serving, cnn_serving, and rnn_serving all have much larger dependency images.
+For this reason, HotSwap is shown to be much faster on cold starts than the Baseline in these cases, by a ratio of 1.2, 1.8, and 2.2.
+Further, there is almost no degradation of performance during warm starts.
+III-E Implementation
+We implement a prototype of HotSwap with AWS Lambda containers and CRIU (Checkpoint and Restore In Userspace) [27].
+HotSwap uses AWS Lambda container templates (Python 3.9) to build serverless functions and collect performance results.
 The dependency loader is embedded into the boot process of an AWS Lambda container.
-When a serverless container receives invocation, it first starts the dependency loader instead of the original “bootstrap” script.
-Then the dependency loader loads dependency images into this container and continually executes the AWS Lambda Runtime Interface [4] and the user code.
-Note that WarmSwap is a provider-side implementation.
-It was necessary to simulate WarmSwap on an AWS EC2 node instead of the online AWS Lambda environment.
+When a serverless container receives an invocation, it first starts the dependency loader instead of the original “bootstrap” script.
+Then the dependency loader loads the dependency image into the container and continually executes the AWS Lambda Runtime Interface [26] and the user code.
+HotSwap is simulated on an AWS EC2 node instead of the online AWS Lambda environment.
+This was necessary, since HotSwap is a provider-side implementation, and it is not possible to modify AWS Lambda.
 Nevertheless, all experiments employ a standard AWS Lambda container and simply execute the container with the EC2 environment.
 The EC2 environment was required because AWS Lambda does not allow the user to control the placement of serverless functions on nodes.
-In particular, it cannot guarantee that a serverless function will be placed on the same node as its associated pool of dependency images.
-The prototype modifies CRIU (Version 3.9) to implement live migration of dependency images. We add two new features supporting the design described in Section 3.2 and Section 3.3.
+The prototype modifies CRIU (Version 3.9) to implement live migration of dependency images. We add two new features supporting the design in Section III-C and Section III-D.
 First, we modified the “criu dump” command to support the dependency pool.
-This ensures that the “criu dump” command will always keep the live dependency image in memory (not killed after migration) and managed by the dependency manager of WarmSwap.
-Additionally, we extend the live dependency image to accept multiple requests.
-Second, we modified the “criu pageserver” command to support the lazy restore policy described in Section 3.3.
-Note that there is a “--lazy page” parameter of the “criu pageserver” command.
+This ensures that the “criu dump” command will always keep the live dependency image in memory (instead of being killed after migration) and managed by the dependency manager of HotSwap.
+Second, we modified the “criu pageserver” command to support the lazy restore policy described in Section III-D.
+We extend the CRIU pageserver to accept multiple requests.
+Note also that there is a “--lazy page” parameter of the “criu pageserver” command.
 However, this parameter migrates all pages upon the execution start.
-We need to implement WarmSwap’s lazy restore policy: loading pages one-by-one.
-Similarly, CRIU’s --lazy page parameter does not fully support WarmSwap’s bulk restore policy.
+We need to implement HotSwap’s lazy restore policy: loading pages one by one.
+Similarly, CRIU’s --lazy page parameter does not fully support HotSwap’s bulk restore policy.
 Some smaller modifications were also required for this policy.
-Finally, it is important that the Python runtime and packages are installed in the same path and with the same version on both the worker node and the container.
-Otherwise, the “criu restore” cannot find the necessary files and reconnect the file descriptors.
-4 Evaluation
-We first present the experimental setup in Section 4.1, including the experimental platform and benchmarks for serverless functions.
-The implementation was described in Section 3.4.
-The evaluation answers the following research questions:
-  • RQ1: How well does WarmSwap improve the cold-start performance of different serverless functions? (Section 4.2)
-  • RQ2: How does each component in WarmSwap contribute to the cold start of different serverless functions? (Section 4.3)
-  • RQ3: How well does live migration in WarmSwap improve the cold-start execution compared with standard migration using simple checkpoint images? (Section 4.4)
-  • RQ4: How do different migration policies (lazy restore and bulk restore) work in WarmSwap? (Section 4.4)
-  • RQ5: How well does WarmSwap perform when sharing dependency images across multiple serverless functions? (Section 4.5)
-4.1 Experimental Setup
-We use a “r5b.large” AWS EC2 machine with two vCPUs, 16 GB memory, and AWS EBS 200 GB SSD, to conduct our experiments.
-The O/S is Amazon Linux with Linux kernel 5.10.199, and the gcc version is 10.5.0.
-As described in Section 3.4, an AWS EC2 node is required to simulate a provider-side WarmSwap implementation, but standard AWS Lambda server containers are used in all experiments.
+We will release our configuration files and prototype code online as open source, to aid in reproducibility.
+IV Evaluation
+We first present the experimental setup in Section IV-A, including the experimental platform and benchmarks for serverless functions. The evaluation answers the following research questions:
+  • RQ1: How well does HotSwap improve the cold start of different serverless functions? (Section IV-B)
+  • RQ2: How does each HotSwap component affect cold starts across serverless functions? (Section IV-C)
+  • RQ3: How well does live migration in HotSwap improve the cold-start execution compared with standard migration using simple checkpoint images? (Section IV-D)
+  • RQ4: How do different migration policies (lazy restore and bulk restore) work in HotSwap? (Section IV-D)
+  • RQ5: How does HotSwap compare to PreBaking [13, 14] when sharing dependency images across multiple serverless functions in a real-world trace? (Section IV-E)
+IV-A Experimental Setup
+We use “r5b.large” AWS EC2 machines with two vCPUs, 16 GB memory, and AWS EBS 200 GB SSD, to conduct our experiments.
+The OS is Amazon Linux with Linux kernel 5.10.199, and the gcc version is 10.5.0.
+As described in Section III-E, an AWS EC2 node is required to simulate a provider-side HotSwap implementation, but standard AWS Lambda server containers are used in all experiments.
 Moreover, all serverless containers are built based on the “public.ecr.aws/lambda/python:3.9” template.
@@
 All experiments use Python-based functions from the FunctionBench [24, 25] serverless benchmark suite.
-Recall that the middleware packages needed by these functions are described in Table 1 in Section 2.3.
-The purposes of these serverless functions are as follows:
-  • helloworld: Minimal function with no dependency packages.
-  • pyase: Text encryption with an AES block-cipher.
-  • chameleon: HTML table rendering.
-  • json_dumps_load: Serialization and de-serialization of JSON.
-  • lr_serving: Serving a logistic regressions model.
-  • cnn_serving: Serving an image classification model based on CNN.
-  • rnn_serving: Serving a sequence generation model based on RNN.
+Recall that the middleware packages needed by these functions and other details are described in Table I in Section II-C.
 Functions needing model parameters (namely, lr _serving, cnn _serving, rnn _serving) download their parameters from AWS S3 when executing cold starts.
-The dependency images created for each serverless function are based on Python and Python packages described in Table 1.
+The dependency images created for each serverless function are based on Python and Python packages described in Table I.
 The cold-start latency discussed in this section refers to the times for: booting at runtime; loading dependencies; and executing the user functions, unless otherwise specified.
 To record the cold-start latency, we first start the serverless container in the worker node and wait for the invocations to eliminate the overhead of container creation.
-4.2 Comparing WarmSwap and Baseline
-To answer RQ1, we evaluate the cold-start latency of WarmSwap and the Baseline with all functions described in Section 4.1.
+IV-B Comparing HotSwap and Baseline
+To answer RQ1, we evaluate the cold-start latency of HotSwap and the Baseline with all functions described in Section IV-A.
 Baseline refers to a traditional AWS Lambda container running offline.
-The Baseline runs the same containers that were built for online AWS Lambda in Section 2.3 on an EC2 machine.
-In this experiment, WarmSwap uses bulk restore (see Section 3.3) because it has the best performance.
+The Baseline runs the same containers that were built for online AWS Lambda in Section II-C on an EC2 machine.
+In this experiment, HotSwap uses bulk restore (see Section III-D) because it has the best performance.
 These serverless containers directly boot the AWS Lambda Runtime Interface during cold starts and use a Python environment and packages pre-installed inside the serverless container.
-(a) Cold Start
-(b) Warm Start
-Figure 5: Comparison of WarmSwap versus Baseline.
-The software dependencies for each function are specified in Table 1 of Section 2.3.
-Note that lr_serving, cnn_serving, and rnn_serving all have much larger dependency images.
-For this reason, WarmSwap is shown to be much faster on cold starts than the Baseline in these cases, by a ratio of 1.2, 1.8, and 2.2.
-Further, there is almost no degradation of performance during warm starts.
-Figure 5a shows that WarmSwap outperforms the Baseline on cold-start latency for serverless functions that require large dependencies.
-WarmSwap is slightly slower than Baseline for the lightweight serverless functions (namely, helloworld, pyase, chameleon, and json_ dumps_load).
-However, WarmSwap significantly outperforms Baseline for serverless functions that need larger packages (i.e., “pandas”, “numpy”, “scikit-learn”, “torch”, and “keras”).
-WarmSwap is 1.2x, 1.8x, and 2.2x faster than Baseline for “lr _serving”, “cnn _serving”, and “rnn _serving”, respectively.
-Figure 5b shows the warm-start latency of WarmSwap and Baseline. For all six serverless functions, WarmSwap has similar warm-start latencies with Baseline.
-The cold start results show that WarmSwap has a performance advantage only on certain serverless functions.
-Specifically, WarmSwap is designed to solve the large dependency-loading problem in the serverless cold start.
+Figure 5a shows that HotSwap outperforms the Baseline on cold-start latency for serverless functions that require large dependencies.
+HotSwap is slightly slower than Baseline for the lightweight serverless functions (namely, helloworld, pyase, chameleon, and json_dumps_load).
+In production, HotSwap would be configured to avoid caching the trivial dependencies associated with these small serverless functions.
+Nevertheless, HotSwap significantly outperforms Baseline for serverless functions that need larger packages (i.e., “pandas”, “NumPy”, “scikit-learn”, “torch”, and “keras”).
+HotSwap is 1.2x, 1.8x, and 2.2x faster than Baseline for “lr _serving”, “cnn _serving”, and “rnn _serving”, respectively.
+Figure 5b shows the warm-start latency of HotSwap and Baseline. For all six serverless functions, HotSwap has similar warm-start latencies with Baseline.
+The cold start results show that HotSwap has a performance advantage only on certain serverless functions.
+Specifically, HotSwap is designed to solve the large dependency-loading problem in the serverless cold start.
 The lightweight functions only import standard and small Python packages.
-So directly booting those functions is faster than using WarmSwap.
-The lightweight functions incur additional overhead because WarmSwap needs to communicate with the dependency manager and migrate the dependency image into the containers, which increases some overhead.
-The warm start results show that WarmSwap’s performance is almost the same as that of the Baseline containers.
-During cold starts, the user code is already executed and necessary pages are already installed.
-WarmSwap does not add overhead to the warm start of the same function.
+So directly booting those functions is faster than using HotSwap.
+The lightweight functions incur additional overhead because HotSwap needs to communicate with the dependency manager and migrate the dependency image into the containers, which increases some overhead.
+The warm start results show that HotSwap’s performance is almost the same as that of the Baseline containers.
+During warm starts, the user code is already executed, and necessary pages are already installed.
+HotSwap does not add overhead to the warm start of the same function.
 Furthermore, there is no difference between the execution times for regular software dependencies and for migrated dependencies.
-4.3 Components of Cold Start Execution Time
-To further analyze WarmSwap when executing a cold start, and answer RQ2, we decompose the cold-start time of WarmSwap and Baseline into three parts.
-Recall the architecture of WarmSwap from Figure 4a.
+IV-C Components of Cold Start Execution Time
+To analyze HotSwap’s cold-start performance and address RQ2, we decompose the cold-start time of HotSwap and Baseline into three parts.
+Recall the architecture of HotSwap from Figure 4a.
 Communication is the process by which the “migration client” communicates with the dependency manager and receives the process metadata.
-The process metadata is a special data structure associated with checkpoints that support CRIU’s lazy migration protocol (see Section 3.2).
-Migration is the process in which the dependency loader executes until the start of user-specific handler functions.
-The definitions of Execution, Boot, and Depedency Initialization are the same as those in Section 2.3.
-Figure 6: Cold-start latency breakdown for WarmSwap (H) and Baseline (B).
-The performance gain of WarmSwap arises because the migration is faster than the software dependency initialization of a Baseline container.
-Figure 6 shows the breakdown of cold-start latency for WarmSwap.
+The process metadata is a special data structure associated with checkpoints that support CRIU’s lazy migration protocol (see Section III-C).
+Migration is the process in which the dependency loader executes until the start of the user-specific handler function.
+The definitions of Execution, Boot, and Dependency Initialization are the same as those in Section II-C.
+Figure 6: Cold-start latency breakdown for HotSwap (H) and Baseline (B).
+The performance gain of HotSwap arises because the migration is faster than the software dependency initialization from scratch.
+Figure 6 shows the breakdown of cold-start latency for HotSwap.
 The communication time is almost constant for all tested serverless functions, ranging from 25 ms to 94 ms according to the sizes of process metadata.
-Table 3 shows the sizes of process metadata for dependency images for each function.
+Table II shows the sizes of process metadata for dependency images for each function.
 The size of process metadata is much smaller than the checkpoint image of the entire dependency.
-Since WarmSwap already migrates most of the memory by the page server, it does not need to transfer much additional information in the process metadata.
-Compared with the dependency-loading time of Baseline, WarmSwap has a faster migration time in the case of functions that have larger dependencies.
-WarmSwap is 2.2x, 3.2x, and 2.5x faster than Baseline for booting dependencies for “lr _serving”, “cnn _serving” and “rnn _serving”, respectively.
-That explains how WarmSwap helps with the cold-start time.
-Since all memory pages are pre-initialized in the dependency image, user functions directly use them after migration.
-Within WarmSwap, the migration time dominates the cold start because migration is the key phase that transfers the dependency image that is needed by the user function.
-4.4 Analysis of Benefits of WarmSwap Components
-This section evaluates the contribution of the different optimization techniques used by WarmSwap in cold starts: live migration (RQ3) and different migration policies (RQ4).
-Since WarmSwap shows performance improvements only for functions with larger dependencies, we focus on the results for “lr_serving,” “cnn_serving,” and “rnn_serving” in this experiment.
-Two migration policies of WarmSwap from Section 3.3 are evaluated:
-WarmSwap Initiative Restore:
-The page server copies all memory pages in parallel after the first “Page Fault.”
-WarmSwap Lazy Restore:
-The page server copies only the accessed pages that trigger a “Page Fault” at a time.
-Two variants of WarmSwap are created as follows:
-WarmSwap without Page Server:
-Prepares a regular checkpoint image for the dependency, copies it to the user container during cold starts, and restores it.
-WarmSwap without Lazy Migration:
-Copies all memory pages through the page server inside the user container before the execution of user code.
-Prototypes Cold Start (s) Warm Start (s)
-lr_serving cnn_serving rnn_serving lr_serving cnn_serving rnn_serving
-WarmSwap Bulk Restore 2.67 2.06 0.89 1.80 0.65 0.004
-WarmSwap Lazy Restore 2.93 3.29 0.76 2.39 2.71 0.004
-w/o Page Server 3.00 2.14 0.91 1.82 0.68 0.004
-w/o Lazy Migration 3.09 2.13 0.89 1.81 0.65 0.004
-Table 2: Comparison of cold-start and warm-start results with: WarmSwap bulk restore, WarmSwap Lazy Restore; WarmSwap without Page Server; and WarmSwap without Lazy Migration.
-Table 2 shows the cold-start and warm-start times for WarmSwap and its variants.
-WarmSwap with bulk restore has the best latency for “lr_serving” and “cnn_serving”, while WarmSwap with lazy restore outperforms the others for “rnn_serving”.
-The cold-start performance difference between bulk restore and lazy restore is related to the number of page faults a function triggers.
-The overhead of lazy restore increases significantly as the number of page faults rises.
-This is because whenever the function triggers a page fault, it needs to pause and wait for the page server to install the specific memory pages.
-For WarmSwap without Page Server, we need to copy a relatively large image to the user container and restore all memory pages before execution.
-The overhead of copying the image is related to the checkpoint sizes, which are listed in Table 3.
-For WarmSwap without Lazy Migration, the cold-start time is slower than WarmSwap because it needs to transfer all the memory pages through the network.
-Note that for “lr _serving”, WarmSwap without Lazy Migration is even slower than WarmSwap without page server.
-This is because “lr _serving” has a smaller checkpoint image size than the other functions.
-All three prototypes (not counting Lazy Copy) have similar warm start times, which demonstrates that those three migration methods do not impact the time for warm starts.
-For Lazy Copy, the performance of warm starts decreases because “lr_serving” and “cnn_serving” trigger extra page faults during the first warm start.
-The performance will become normal for the following warm starts.
-Based on these experiments, we believe bulk restore is more suitable in practice since the provider usually does not know how many page faults the target program will trigger.
-Moreover, it has no impact on warm-start performance.
-Functions Metadata Size (MB) Image Size (MB)
+TABLE II: The size of data that needs to be transferred from the worker node to the function container of HotSwap and HotSwap without Page Server.
+Functions Metadata (MB) Image (MB)
 helloworld 0.88 8.1
@@
 rnn_serving 12 200
-Table 3: The size of data that needs to be transferred from the worker node to the function container of WarmSwap and WarmSwap without Page Server.
-4.5 Case Study: Dependency Image Sharing Across Serverless Functions
-In this section, we conduct simulation experiments to show the benefit of WarmSwap on sharing dependency images across different serverless functions to answer RQ5.
-We modify the “rnn_serving” functions by adding new code (using NumPy and torch) to the original user handler to create 10 different serverless functions.
-Assuming these 10 serverless functions are less frequently invoked, there is at most one instance for each function.
-However, the provider would still wish to gain the benefits of warm start even on these functions, if the cost is reasonable, in order to reduce the cold-start overhead.
-Figure 7 compares WarmSwap and a checkpoint image-based method, Prebaking [23].
-We implement Prebaking following the description in their paper, by checkpointing the entire previous warm serverless container with CRIU [7] to a function image and restoring it during the cold start.
-Note that, unlike WarmSwap and Baseline, the cost of container creation is larger in the case of Prebaking.
-Therefore, Figure 7 also includes the overhead of container creation for both methods.
-Figure 7: Comparison of the cold-start latency and required cache/memory of WarmSwap and Prebaking Image.
-The experiment is based on 10 different serverless functions that require the same dependency image (Python+numpy+torch).
-Figure 7 shows the accumulated cold-start latency (the sum of average cold-start latency for the 10 functions), and the required cache in memory needed by the optimization.
-WarmSwap significantly outperforms Prebaking in the cold-start performance (4.8 times faster), since restoring a large container with CRIU is much slower than migrating only the dependency.
-Furthermore, the provider only needs 260 MB memory to warm up the 10 different functions when using WarmSwap.
-In Prebaking, the provider needs to keep 10 checkpoint images (about 1.8 GB) in the cache for the 10 different functions.
-Hence, warming up the less frequently used functions with WarmSwap is cheaper than Prebaking.
+Compared with the dependency-loading time of Baseline, HotSwap has a faster migration time in the case of functions that have larger dependencies.
+HotSwap is 2.2x, 3.2x, and 2.5x faster than Baseline for booting dependencies for “lr _serving”, “cnn _serving” and “rnn _serving”, respectively.
+That explains how HotSwap helps with the cold-start time.
+Since all memory pages are pre-initialized in the dependency image, user functions directly use them after migration.
+Within HotSwap, the migration time dominates the cold start because migration is the key phase that transfers the dependency image that is needed by the user function.
+IV-D Ablation Study
+This section evaluates the contribution of the different optimization techniques used by HotSwap in cold starts: live migration (RQ3); and different migration policies (RQ4).
+Since HotSwap shows performance improvements only for functions with larger dependencies, we focus on the results for “lr_serving,” “cnn_serving,” and “rnn_serving”.
+TABLE III: Comparison of cold-start and warm-start results with: HotSwap bulk restore, HotSwap Lazy Restore; HotSwap without Page Server; and HotSwap without Lazy Migration.
+Prototypes Cold Start (s) Warm Start (s)
+lr_serving cnn_serving rnn_serving lr_serving cnn_serving rnn_serving
+HotSwap Bulk Restore 2.67 2.06 0.89 1.80 0.65 0.004
+HotSwap Lazy Restore 2.93 3.29 0.76 2.39 2.71 0.004
+w/o Page Server 3.00 2.14 0.91 1.82 0.68 0.004
+w/o Lazy Migration 3.09 2.13 0.89 1.81 0.65 0.004
+Two migration policies of HotSwap from Section III-D are evaluated:
+HotSwap Bulk Restore:
+The page server copies all memory pages in parallel after the first “page fault”.
+HotSwap Lazy Restore:
+The page server copies only the accessed pages that trigger a “page fault”.
+Two variants of HotSwap are created as follows:
+HotSwap without Page Server:
+Prepares a regular checkpoint image for the dependency, copies it to the user container during cold starts, and restores it.
+HotSwap without Lazy Migration:
+Copies all memory pages through the page server inside the user container before the execution of user code.
+Table III shows the cold-start and warm-start latencies for HotSwap and its variants.
+HotSwap with bulk restore has the best latency for “lr_serving” and “cnn_serving”, while HotSwap with lazy restore outperforms the others for “rnn_serving”.
+The cold-start performance gap between bulk and lazy restore depends on page faults, as each fault in lazy restore pauses execution to wait for the page server.
+For HotSwap without a Page Server, we copy a relatively large image to the user container and restore all memory pages before execution.
+The overhead of copying is related to the checkpoint sizes listed in Table II.
+For HotSwap without Lazy Migration, the cold-start time is slower than HotSwap because it needs to transfer all the memory pages through the network before the start of function execution.
+Note that for “lr _serving”, HotSwap without Lazy Migration is even slower than HotSwap without page server.
+This is because “lr _serving” has a smaller checkpoint image size than the other functions.
+All three prototypes (excluding Lazy Restore) have similar warm-start times, demonstrating that these migration methods have negligible impact on warm-start latency.
+For Lazy Restore, the performance of warm starts decreases because “lr_serving” and “cnn_serving” trigger extra page faults during the first warm start.
+Based on these experiments, we believe bulk restore is more suitable in practice since the provider usually does not know how many page faults the target program will trigger.
+IV-E Case Study: Dependency Image Sharing Across Serverless Functions
+In this section, we conduct simulation experiments to demonstrate the benefits of HotSwap in sharing dependency images across different serverless functions, addressing RQ5.
+We use real-world serverless function traces from Azure [11], previously discussed in Section II-B of this paper, as the basis for our tests.
+We modified the “rnn_serving” function by adding new code (utilizing NumPy and Torch) to create 10 different serverless functions.
+For each function, we assigned random traces and measured latency over a two-week period.
+The keep-alive time was set to $15$ minutes.
+Since HotSwap is designed to optimize performance for less frequently used functions, we tested it under various invocation frequencies.
+This experiment compares HotSwap with a checkpoint image-based method, Prebaking [13, 14].
+Prebaking is chosen as a well-known system that employs CRIU-based checkpoint/restore to accelerate serverless cold starts.
+We implemented Prebaking as described in their paper, by checkpointing the entire previously warm serverless container using CRIU [27] to create a function image, which is then restored during a cold start.
+To enhance fairness, we store all prebaking images in memory.
+Note that the cold start latency measured in this experiment includes the container creation overhead for both methods.
+Figure 7: Comparison of average latency and required cache/memory between HotSwap and Prebaking[13, 14] under Azure Function traces [11].
+The experiment involves 10 different serverless functions, all using the same dependency image ( NumPy + Torch).
+Each function is assigned a real-world invocation trace from Azure.
+The left figure shows latency across different invocation rate groups, ranging from the lowest to the highest quartile.
+Figure 7 shows the average latency and required cache for HotSwap and Prebaking.
+The average latency is calculated from all invocations of the ten different functions.
+To evaluate different invocation rate scenarios, we categorized the Azure traces into quartiles based on invocation rate.
+The “Lowest Quartile” represents experiments on function invocation traces with rates below the 25th percentile, while the other three groups correspond to “25-50%”, “50-75%”, and above the 75th percentile, respectively.
+The latency results indicate that HotSwap outperforms Prebaking by 1.4x across all trace groups.
+This improvement is due to HotSwap’s superior cold-start performance.
+The average latency decreases as the invocation rate increases, resulting in more warm starts.
+Notably, the 75th percentile of invocation rates from Azure traces [11] is 0.04 calls per minute.
+This demonstrates that the majority of serverless functions are, in fact, used infrequently.
+Furthermore, the provider requires only 260 MB of memory to warm up the 10 different functions when using HotSwap.
+In contrast, Prebaking requires the provider to maintain 10 checkpoint images (approximately 2.4 GB) in the cache for the same functions.
+Therefore, warming up the less frequently used functions with HotSwap is more cost-effective than Prebaking.
+This benefit can increase further if there are more functions that share the same dependency image.
 Although function-specific tuning methods, like vHive [17] and Catalyzer [15], achieve faster container restore, the provider must create and keep all function images in cache to optimize the less frequently used serverless functions.
-Thus, WarmSwap provides a practical method for providers to optimize cold starts for all functions at a reasonable cost.
-5 Related Work
-5.1 Function-Specific Tuning
-One straightforward way to perform function-specific tuning is by using the Checkpoint/Restore techniques like CRIU [7].
-One can create a checkpoint image of a specific warm serverless function and restore that image instead of initializing it from scratch during a cold start.
-For example, Prebaking [23] utilized transparent checkpoint technology (CRIU) to locally checkpoint the serverless function.
-This approach involves restarting the process during cold starts.
-Catalyzer, as presented by Du et al.
-[8], confronts the intricate task of swiftly restoring serverless sandboxes, containers, virtual machines, and more.
-This innovative approach adeptly reinstates a virtualization-based function instance from a structured checkpoint image.
-Moreover, Catalyzer advances this restoration process by enabling on-demand recovery of both user-level memory states and system states.
-Notably, Catalyzer introduces the concepts of “sandbox fork” and “fork boot” to effectively tackle the challenge of warm starts, particularly when one or more hot instances already exist.
-Building on checkpoint images, researchers have proposed more advanced optimizations.
-The authors of Catalyzer extended their contributions through MITOSIS [26].
-This work takes the notion of sandbox forking further by leveraging Remote Direct Memory Access (RDMA) for remote forking.
-This innovation enables rapid warm starts for multiple sandboxes and supports stateful serverless computing.
-This extension enhances the versatility and applicability of the sandbox forking concept, facilitating efficient warm start strategies and accommodating stateful serverless scenarios.
-Building upon the foundational work of Catalyzer, another study[25] introduces a noteworthy approach termed “hot page record-and-prefetch technology” aimed at expediting sandbox restoration and serverless function execution.
-The underlying concept revolves around recognizing that not all data within the C/R image will be utilized following restoration.
-Additionally, the SEUSS framework [6] presents a unique C/R-like solution.
-In contrast to other approaches, SEUSS focuses on the restoration of function unikernel snapshots, sidestepping resource-intensive initialization steps.
-An intriguing aspect of SEUSS is its incorporation of page-level sharing across the entire software stack required for function execution.
-SEUSS encompasses not only unikernel binaries but also runtime snapshots and user code snapshots.
-There are more ways to conduct function-specific tuning to speed up cold starts by analyzing user function code.
-Fasslight [17] optimizes code loading latency by loading only indispensable code.
-To implement this, Fasslight constructs the function-level call graph, separates other code, and loads only indispensable code to reduce cold start overhead.
-Function-specific tuning usually achieves extremely high-speed cold starts, as reported by MITOSIS [26] and vHive [25].
-However, as we analyzed in Section 2, these methods are not suitable for all serverless functions in production, especially those that are less frequently used.
-Additionally, some methods are difficult to implement on the cloud provider side because they require the provider to record [25] or analyze user code [17] or containers.
-5.2 Infrastructure Optimization
-For infrastructure optimization, the easiest way is to use a better hardware infrastructure.
-Further, there are plenty of system designs that help faster loading and initializing the functions by virtualization and caching systems.
-Firecracker [1] by AWS Lambda is a Virtual Machine Monitor that utilizes the Linux Kernel’s KVM virtualization infrastructure to offer lightweight virtual machines.
-Firecracker significantly reduces the boot time of serverless instances and enhances switching performance.
-AWS further reports its container optimization, which has implemented sharing across serverless functions in its approach to loading containers on-demand [5].
-This is achieved through transforming serverless containers into data chunks, which are then deduplicated [18, 27] and shared across functions.
-These data chunks are efficiently stored within a three-tiered cache system.
-The primary focus of this work is on optimizing storage and facilitating the swift movement of larger serverless containers, all aimed at accelerating cold-start times.
-The inherent nature of serverless functions has opened avenues for sharing not only within the same function but also across different functions.
-This is particularly valuable for frequently utilized functions that may accumulate millions of instances.
-Developers often opt for identical containers, runtimes, or even packages.
-In a recent report, AWS revealed that a significant 80% of newly uploaded serverless containers lack new data chunks [5].
-Building on this insight, researchers extend infrastructure optimization through shareable resources.
-Pagurus [14], for instance, explores container-level sharing.
-Recognizing that certain user functions extensively share software packages, this approach allows new functions to borrow containers from existing ones, thereby reducing cold-start overhead.
-However, it’s crucial to acknowledge that sharing entire containers could introduce additional security challenges.
-HotC [24] focuses on reusing container runtimes to minimize container initialization overhead.
-HotC maintains a live container runtime pool, analyzes user input or configuration files, and provides available runtimes for immediate reuse.
-In this work, we propose a feasible solution to extend infrastructure optimization for reducing dependency loading overhead.
-5.3 Dependency Loading Overhead
-Several previous works have recognized the importance of software dependency loading overhead.
-The SOCK framework, as explored in Oakes et al.’s work [21], analyzed the impact of loading Python packages on overhead.
-Their findings revealed that incorporating numerous popular libraries during the import process adds approximately 100 milliseconds to the cold-start time.
-Consequently, they devised an innovative strategy involving the creation of a lean container and a three-tiered Python library caching system.
-This approach effectively accelerates cold-start times.
-Similarly, the FlashCube framework [16] addresses the delays caused by initializing function runtimes.
-To tackle this issue, FlashCube introduces a novel container framework.
-Instead of constructing a container from scratch, FlashCube assembles it by combining a selection of pre-existing general container components, including namespaces, cgroups, and runtimes.
-This approach maintains a template runtime process for each programming language (such as Python, Java, and Node.js).
-Photons [9] also recognize language runtime and library redundancy.
-They co-locate multiple instances of the same function within the same runtime, allowing concurrent invocations to share dependencies and application states.
-Note that existing solutions attempt to pre-warm dependencies at the container level, requiring providers to pre-warm containers or container components for serverless functions.
-In contrast, WarmSwap is a lightweight solution that can be easily deployed across multiple nodes while effectively addressing the dependency loading problem.
-6 Conclusion
-In conclusion, this paper presents WarmSwap, a provider-side solution for accelerating dependency loading in serverless cold starts in production serverless platforms.
-WarmSwap is a complementary work that extends infrastructure optimization in serverless environments by representing software dependencies as middleware.
-It is designed to benefit the many less frequently used serverless functions, which is often not practical in a function-specific approach.
-WarmSwap is primarily beneficial when the dependencies of the serverless functions are large and when they incur significant latency during initialization.
-Containers with short initialization times already perform well during cold start, but recent years have seen greater use of complex software dependencies like AI/ML.
-It is in this regime that WarmSwap is expected to be used.
-By using WarmSwap, a provider can implement a single pool of dependency images across its many computer hosts.
-This relieves the user of additional work.
-Since dependency images represent public and commonly used middleware, they can be shared among a whole category of serverless functions.
-Unlike previous research, the provider only needs to pre-warm and migrate the dependency process itself instead of entire containers.
-We implemented a prototype of WarmSwap based on offline AWS Lambda containers, and the evaluation shows that WarmSwap reduces the dependency loading overhead and warm-up cost across many serverless functions.
+Thus, HotSwap provides a practical method for providers to optimize cold starts for skewed workloads at a reasonable cost.
+V Related Work
+Pre-Warm.
+One straightforward way to optimize cold starts is to prevent them from happening.
+Prior research has explored various approaches such as adjusting keep-alive policies, leveraging application-level composition [28], mining invocation patterns [9], and applying machine learning models [29, 30, 31], in order to predict function usage and proactively pre-warm containers [10, 8, 11].
+However, these strategies often lead to resource inefficiency due to inaccurate predictions, especially in dynamic production environments.
+HotSwap takes a different approach by directly reducing cold-start overhead instead of relying on speculative pre-warming.
+Function-Specific Optimization.
+A common method for reducing cold-start latency is leveraging Checkpoint/Restore (C/R) techniques such as CRIU [27].
+Prebaking [13, 14] applies CRIU to checkpoint entire serverless functions, significantly speeding up start-up across runtimes.
+Catalyzer [15] improves sandbox restoration by efficiently resuming serverless functions across sandboxes.
+MITOSIS [18] extends this approach by introducing remote sandbox forking using RDMA.
+Further extending this idea, Ustiugov et al.
+[17] propose a “hot page record-and-prefetch” technique to accelerate sandbox restoration and function execution.
+SEUSS [16] takes a unikernel-based approach, restoring function snapshots while avoiding resource-heavy initialization.
+Fasslight [32] optimizes code loading latency by loading only essential code.
+Although effective for high-frequency serverless functions, these methods may not suit skewed workloads in production, where checkpointing and analysis overhead can outweigh the benefits.
+Furthermore, a recent user-perspective study [33] suggests that Node.js performs better than Java for infrequently invoked functions.
+However, providers have no control over which language users choose to implement their functions.
+Cross-Function Optimization and Dependency Loading.
+Cross-function optimizations improve performance by sharing resources across functions.
+Pagurus [34] reduces cold-start overhead via container reuse, while HotC [35] maintains a pool of preloaded runtimes for reuse based on function inputs.
+Software dependency loading is a key factor in cold-start latency.
+SOCK [19] uses pre-initialized Zygote containers and multi-tier caching to speed up Python library loading.
+(In fact, SOCK uses a generalization of the original Android-based Zygotes.) FlashCube [21] builds containers from prebuilt components to reduce initialization time.
+Photons [36] co-locates function instances in a shared runtime to enable dependency reuse.
+TrEnv [20] leverages remote memory pools via CXL/RDMA to share execution environments.
+SLIMSTART [7] removes infrequently used libraries through runtime profiling and automated transformation.
+HotSwap aims to introduce a more lightweight and easily implementable solution to reduce dependency-related overhead.
+Infrastructure and Storage Systems.
+Infrastructure optimization can further reduce cold-start latency by improving storage and caching mechanisms.
+Firecracker [23], a micro-VM developed by AWS Lambda, leverages the Linux Kernel’s KVM virtualization to provide lightweight virtual machines, significantly reducing boot times and improving context-switching performance.
+AWS also optimizes serverless container sharing by deduplicating data chunks across functions, revealing that 80% of uploaded containers contain redundant data [22, 37, 38].
+These infrastructure optimizations are complementary to HotSwap, as they focus on different aspects of serverless performance improvements.
+VI Limitations
+HotSwap is specifically designed for to be active during the dependency loading phase.
+HotSwap targets the overhead associated with dependency loading and is most effective when this overhead is substantial.
+For functions with only trivial dependencies, using HotSwap to accelerate dependency loading may not be necessary.
+From the provider’s perspective, this also means that it is not required to cache all dependencies, only those that are popular and large.
+Evaluation on Other Runtimes and Languages.
+In this paper, we evaluate HotSwap only on the Python runtime.
+We chose Python because it is one of the most popular languages for serverless functions, particularly in data processing, automation, and machine learning tasks.
+Moreover, many widely used Python libraries are large and contribute significantly to cold start overhead.
+However, since HotSwap relies on process-level migration, it could, in theory, be applied to other runtimes and languages as well—especially in scenarios where dependency loading overhead is a major concern.
+Centralized Dependency Management.
+The Dependency Manager introduced in Section III-B is centralized in our prototype design and may become a bottleneck when handling numerous migration requests.
+In practice, cloud providers can implement HotSwap as a distributed caching system with multiple Dependency Managers operating within a cluster.
+VII Conclusion
+In conclusion, this paper presents HotSwap, a cross-function optimization designed to accelerate dependency loading in serverless cold starts.
+HotSwap provides a strong benefit in cases of skewed workloads, with many infrequently used serverless functions (the problem of a “long tail” in the distribution).
+It primarily optimizes the dependency loading phase, a major bottleneck in current production clouds.
+By enabling a shared pool of dependency images across multiple compute hosts, HotSwap allows providers to efficiently manage commonly used middleware.
+We implemented a HotSwap prototype based on offline AWS Lambda containers, and our evaluation demonstrates that it significantly reduces dependency loading overhead and warm-up costs across various serverless functions.
+Acknowledgment. We thank our shepherd, Yehia Elkhatib, and the anonymous reviewers for their constructive feedback. This work is supported in part by NSF Awards 1910601 and 2124897.